### v2

# Alternating Direction Method of Multipliers AR AutoRally

ADMM
Alternating Direction Method of Multipliers
AR
AutoRally
CEM
Cross-Entropy Method
CNN
Convolutional Neural Network
ConvLSTM
Convolutional Long Short Term Memory
CRTP
Curiously Recurring Template Pattern
CTG
Cost-To-Go
CVaR
Conditional Value-at-Risk
DDP
Differential Dynamic Programming
DMD-MPC
Dynamic Mirror Descent Model Predictive Control
EKF
Extended Kalman Filter
EM
Expectation Maximization
ES
Evolutionary Strategies
GMM
Gaussian Mixture Model
GP
Gaussian Process
GRU
Gated Recurrent Unit
FT
Fourier Transform
FCNN
Fully-Connected Neural Network
iCEM
improved Cross-Entropy Method
iFFT
inverse Fast Fourier Transform
iLQR
iterative Linear Quadratic Regulator
IS
Importance Sampler
KL
Kullback Leibler
LSTM
Long Short Term Memory
MOO
Multi-Objective Optimization
MBRL
Model-Based Reinforcement Learning
MPC
Model Predictive Control
MPPI
Model Predictive Path Integral
NLN
Normal log-Normal
NN
Neural Network
PCA
Principle Component Analysis
PDF
Probability Density Function
PSD
Power Spectral Density
RCNN
Region-based Convolutional Neural Network
RL
Reinforcement Learning
ROS
Robot Operating System
RMPPI
Robust-MPPI
RPN
Region Proposal Network
RRT
Rapidly Exploring Random Tree
SQP
Sequential Quadratic Programming
VaR
Value-at-Risk
YOLO
You Only Look Once
MPPI-Generic: A CUDA Library for Stochastic Trajectory Optimization
Bogdan Vlahov1, Jason Gibson1, Manan Gandhi1, Evangelos A. Theodorou1 1 Autonomous Control and Decision Systems Lab, Georgia Institute of Technology, Atlanta GA 30313 USA (Corresponding author: Bogdan Vlahov)
(March 10, 2025)
Abstract
This paper introduces a new C++/CUDA library for GPU-accelerated stochastic trajectory optimization called MPPI-Generic. It provides implementations of Model Predictive Path Integral control, Tube-Model Predictive Path Integral Control, and Robust Model Predictive Path Integral Control, and allows for these algorithms to be used across many pre-existing dynamics models and cost functions. Furthermore, researchers can create their own dynamics models or cost functions following our API definitions without needing to change the actual Model Predictive Path Integral Control code. Finally, we compare computational performance to other popular implementations of Model Predictive Path Integral Control over a variety of GPUs to show the real-time capabilities our library can allow for. Our library can be found at: https://acdslab.github.io/mppi-generic-website/
I Introduction
As robotics and autonomy continue to grow, the choice of algorithms and methods used to plan and control complex systems in these fields start to become more selective. These algorithms need to be able to handle the potentially intricate dynamics found in robotic systems and allow for complex cost function representations. They also must be responsive to new data as environments and systems change. Finally, the planning and control methods used should run in real-time so that the robot can continue to perform without unnecessary pauses. We can see these requirements show themselves in a self-driving car example. The dynamics need to be aware of changing road conditions while the cost function has to capture multiple goals such as avoiding other cars, following traffic laws, and getting to the destination. The planner has to be fast enough to react to other drivers stopping suddenly or debris on the road.
The approaches to Model Predictive Control (MPC) optimization can be delineated into two types of methods: gradient-based and sampling-based. Gradient-based methods such as iterative Linear Quadratic Regulator (iLQR) [1], Differential Dynamic Programming (DDP) [2] and Sequential Quadratic Programming (SQP) [3] generally rely on restrictions to the dynamics and cost functions such as being continuously differentiable. But in exchange for those restrictions, they can produce controls that minimize the cost function in a computationally-efficient manner. Sampling-based methods, such as Model Predictive Path Integral (MPPI) [4] or Cross-Entropy Method (CEM) [5], can relax these requirements and allow for arbitrary functions but come at the cost of requiring many samples to properly estimate the optimal control. These arbitrary functions remove the need for convexification or smoothing which can cause unnecessary conservatism. One way to address this computational expense is to push the computation out of the CPU and onto a GPU, where the parallelization of sampling can be better utilized. By taking advantage of the GPU, we make sampling-based methods usable in real-time and also provide enough samples to get optimal solutions.
First introduced in [4], MPPI is a stochastic MPC algorithm derived using the information theoretic dualities between relative entropy and free energy. Experiments in [4] included off-road navigation using the GT-Autorally vehicle [6] and demonstrated for the first time utilization of GPU for sampling-based MPC on real hardware. Follow-up works included the derivation of MPPI for the case of non-affine dynamics [7, 8] and extensions to multi-layer control architectures that incorporate the benefits of iLQR to increase robustness to disturbances and model errors. These extensions include the Tube-based MPPI [9], Robust-MPPI (RMPPI) [10, 11] architectures which consists of two layers of control. Further discussion of extensions to MPPI is left to Section II-B.
Besides the obvious use-case of MPC, alternative use-cases of MPPI can be found in the Model-Based Reinforcement Learning (MBRL) literature [12, 13, 14, 15, 16, 17]. In MBRL, MPPI is used to seek data from a simulation environment to learn the underlying value function and corresponding policy. Finally, while the majority of prior work on MPPI has direct applications to planning and control for systems in robotics and autonomy such as quad-rotors [18, 19], terrestrial [4, 20, 21, 22, 23], sea surface [24, 25] and underwater [26] vehicles, manipulators [27], and systems with multi-body dynamics [28], there are notable applications to other domains of science and engineering. These include control of HVAC systems [29, 30], chemical reactors [31], and pulse width modulation rectifiers [32].
Given the far reaches of the MPPI algorithm, it is important to provide a solid and flexible computational framework from which researchers can make use of previous advancements to push their own work forwards. That framework needs to not only provide many built-in options to allow for testing on different platforms but also allow for researchers to develop new ideas on top of. Finally, that library needs to be able to support researchers all the way to hardware deployment with the ability to run in real-time even on older hardware platforms.
With this inspiration in mind, we introduce our controls and planning optimization library, MPPI-Generic. It is written in C++/CUDA and contains multiple dynamics and cost functions to allow for researchers to begin using them in complex robotics scenarios. In addition, it also allows for researchers to create their own dynamics or cost functions that take advantage of the GPU-accelerated controllers. It provides implementations of MPPI[4], Tube-MPPI[9], and RMPPI [11] as well as an API for implementations of new sampling-based controller. Initial versions of this library have already been used in a variety of hardware and software experiments [20, 24, 11, 33, 34]. To the best of our knowledge, this is the first MPPI library implementation to provide GPU-acceleration with real-time performance, multiple existing dynamics and cost functions, replaceable sampling distributions and controller algorithms, and extensibility options for researchers to create new components.
Our contributions are summarized as follows:
  • We provide a NVIDIA GPU-accelerated Library [^1] Library code is located at https://github.com/ACDSLab/MPPI-Generic for MPPI.
  • We provide a C++/CUDA API that allows for arbitrary dynamics and cost function definitions while maintaining high computational performance. We show simple examples of extending various parts of the library to give researchers an idea of how they could customize this library for their own needs.
  • We show computational comparisons with other popular MPPI libraries across a variety of computational hardware.
The rest of the paper is organized as follows: In Section II, we provide the general problem formulation of a stochastic trajectory optimization problem, introduce various extensions to MPPI, and discuss the building blocks of the MPPI-Generic Library. In Section III, we discuss implementation details of our library and performance parameters available to the user to tweak. In Section IV, we show how to use the library through coding examples. In Section V, we show computational comparisons of our library against other implementations of MPPI. Finally, we state our conclusions in Section VI.
II Background on Stochastic Trajectory Optimization
Consider a general nonlinear system with discrete dynamics and cost function of the following form:
${\bf x}_{t+1}$ $={\bf F}\left({\bf x}_{t},{\bf u}_{t}+{\bf n}_{t}\right)$ (1)
$\operatorname{\bf J}(X,U)$ $=\phi({\bf x}_{T})+\sum_{t=0}^{T-1}{\bf\ell}\left({\bf x}_{t},{\bf u}_{t}\right),$ (2)
where ${\bf x}\in\mathbb{R}^{n_{x}}$ is the state, ${\bf u}\in\mathbb{R}^{n_{u}}$ is the control, ${\bf n}$ is assumed to be zero-mean Gaussian disturbances with variance $\Sigma$ in the control channel, $T$ is the time horizon, $X$ is a state trajectory $[{\bf x}_{0},{\bf x}_{1},...,{\bf x}_{T}]$, $U$ is a control trajectory $[{\bf u}_{0},{\bf u}_{1},...,{\bf u}_{T-1}]$, $\phi$ is the terminal cost, and ${\bf\ell}$ is the running cost. It is important to note that existing MPPI-Generic dynamics assume Euler integration by default but can be modified if desired by developing a custom dynamics class.
II-A The MPPI Algorithm
While MPPI was originally derived from a path-integral approach, there have been other derivations that lead to slightly different update rules. We will briefly go over the information-theoretic derivation [8] to underpin the specific update rule we use in MPPI-Generic. We also briefly discuss other variations of MPPI that are included in this library.
II-A1 Information Theoretic Derivation
We shall follow the information-theoretic derivation from [8] with some changes. In [8], the running cost definition was only based on the state trajectory and the authors showed how a control cost would emerge naturally from the importance sampling scheme. However, the derivation does not require this limited running cost definition, so we generalize our running cost to $\ell({\bf x}_{t},{\bf u}_{t})$. We use the following shorthand to denote the cost for a control trajectory $U$ applied to the starting state, ${\bf x}_{0}$,
$\begin{split}\mathcal{J}\left(U\right)&:={\bf J}(X,U)\\ \text{s.t. }&{\bf x}_{t+1}={\bf F}\left({\bf x}_{t},{\bf u}_{t}\right)\end{split}$ (3)
Our goal is to find a control trajectory, $U^{*}$, that minimizes Eq. 2 through the use of sampling. We start by defining the free energy of our system as
$\mathcal{F}\left(\mathcal{J},\mathbb{P},\lambda;{\bf x}_{0}\right)$ $=-\lambda\ln\left(\mathbb{E}_{V\sim\mathbb{P}}\left[\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right)\right]\right)$ (4)
where ${\bf x}_{0}$ is the initial state, $\lambda\in\mathbb{R}^{+}$ is the inverse temperature, and $V$ is the control trajectory sampled from some base distribution $\mathbb{P}$ with Probability Density Function (PDF) $p$. The base distribution $\mathbb{P}$ can be any distribution, making it potentially computationally intractable. Introducing a new distribution $\mathbb{Q}$ that is absolutely continuous with $\mathbb{P}$, we use importance sampling and Jensen’s Inequality to get the following,
$\mathcal{F}\left(\mathcal{J},\mathbb{P},\lambda;{\bf x}_{0}\right)$ $\leq\mathbb{E}_{V\sim\mathbb{Q}}\left[\mathcal{J}\left(V\right)\right]+\lambda\mathbb{KL}\left(\mathbb{Q}\ ||\ \mathbb{P}\right)$ (5)
where $\mathbb{KL}\left(\cdot\ ||\ \cdot\right)$ is the Kullback Leibler (KL) divergence between $\mathbb{Q}$ and $\mathbb{P}$. This introduces an upper bound on our free energy when we sample from $\mathbb{Q}$ instead of the original $\mathbb{P}$; this upper bound becomes a strict equality if $\mathbb{Q}^{*}$ has the following density:
$q^{*}(V)$ $=\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right)p(V)$ (6)
$\eta$ $=\mathbb{E}_{\hat{V}\sim\mathbb{P}}\left[\exp\left(-\frac{1}{\lambda}\mathcal{J}(\hat{V})\right)\right]$ (7)
While $\mathbb{Q}^{*}$ is an optimal distribution that minimizes the free energy bound, the optimal PDF, $q^{*}(V)$, still relies on the PDF of the original distribution $\mathbb{P}$, meaning it still might be computationally intractable. Thus, we introduce a third distribution $\mathbb{S}_{\theta}$ that has controllable parameters $\theta$ and minimizes the KL divergence between $\mathbb{S}_{\theta}$ and $\mathbb{Q}^{*}$,
$\theta^{*}$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{KL}\left(\mathbb{Q}^{*}\ ||\ \mathbb{S}_{\theta}\right)$ (8)
$=\operatorname*{arg\,max}_{\theta\in\Theta}\mathbb{E}_{V\sim\mathbb{Q}^{*}}\left[\ln\left(s\left(V|\theta\right)\right)\right].$ (9)
Eq. 9 gives us an update rule for any probability distribution with parameters to match the optimal distribution. Leaving this aside for the moment, we show how to calculate the optimal control sequence by sampling from $\mathbb{S}_{\theta}$. From there, we will show how the update rule and the optimal control sequence can relate to each other depending on the choice of distributional family, $\mathbb{S}_{\theta}$. For the optimal control to use at time $t$, we shall look to the mean of the optimal distribution $\mathbb{Q}^{*}$,
${\bf u}_{t}^{*}$ $=\mathbb{E}_{V\sim\mathbb{Q}^{*}}\left[{\bf v}_{t}\right]$ (10)
[…]
Using importance sampling to sample from $\mathbb{S}_{\theta}$ as well as Eq. 6 gives the following,
[…]
$=\mathbb{E}_{V\sim\mathbb{S}_{\theta}}\left[w(V){\bf v}_{t}\right]$ (13)
$w(V)$ $=\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right)\frac{p\left(V\right)}{s\left(V|\theta\right)}$ (14)
The only term in $w(V)$ that is unaccounted for is the importance sampling weight of $\mathbb{P}$ and $\mathbb{S}_{\theta}$.
In [8], the base and parameterized distributions were both assumed to be Gaussian distributions with constant variance across time, $\Sigma$. The mean trajectory of $\mathbb{P}$ was assumed to be ${\bf 0}_{n_{u}\times T}$ while the mean trajectory of $\mathbb{S}_{\theta}$ was the parameter to optimize, i.e. $\theta={\bf U}^{*}=\left\{{\bf u}^{*}_{t}\right\}_{t=0}^{T-1}$. This led to the specific form of
$w(V)$ $=\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\mathcal{J}(V)-\sum_{t=0}^{T-1}\big{(}{\bf v}_{t}-\frac{1}{2}{\bf u}^{*}_{t}\big{)}^{\top}\Sigma^{-1}{\bf u}^{*}_{t}\right)$ (15)
and also meant that the solutions to Eqs. 9 and 13 were the same.
II-B Other MPPI Modifications
Beyond the derivations of MPPI, there have been other modifications that address various limiations of MPPI. A Tube-based MPPI controller [9] was created in order to improve robustness to state disturbances. It made use of a tracking controller to track the real system back to a nominal system that ignored state disturbances resulting in large costs. The nominal system would use an initial state produced from the dynamics equation in Eq. 1 for a new iteration of MPPI instead of the real system’s state estimate. Both the real and the nominal trajectories are calculated using MPPI while the tracking controller was iLQR. In this setup, the tracking controller would always be the one sending controls to the system and as MPPI was not aware of the tracking controller, it could end up fighting against the tracking controller. In order to address that, RMPPI was developed in [10, 11], which applied the tracking controller feedback within the samples MPPI used. RMPPI also chose the initial state for the nominal system using a constrained optimization problem that tried to keep the nominal state as close to the real state without causing the resulting trajectory to have a cost larger than a given cost threshold $\alpha$. Choosing the nominal state in this way ensured there was an upper bound on how quickly the optimal trajectory’s cost could grow due to disturbance. However, this improved robustness to state disturbances can struggle when the cost function itself changes over time due to new information. Our library contains implementations of these algorithmic improvements to MPPI as different controllers are the best choice in different scenarios.
Since this first wave of publications on MPPI, there has been a plethora of extensions and improvements. These include sampling efficiency improvements such as Covariance Control MPPI [35], Colored MPPI [20], Residual MPPI [36], Biased-MPPI [37], Spline-Interpolated Model Predictive MPPI [19], Stein Variational Guided MPPI [38], log-MPPI [21], CoVO-MPC [39], U-MPPI [40], o-MPPI [41], and Smooth MPPI [42]; there have also been robustness improvements such as Tsallis-MPPI[43], Constrained Covariance Steering MPPI [44], Multi-Modal MPPI [45], and Risk-Aware MPPI [46]. Some of the sampling efficiency methods have also been implemented, shown in Section IV-C3, but the rapid development of the community makes it nearly impossible to keep up.
II-C MPPI-Generic’s MPPI Implementation
In practice, MPPI-Generic uses the update rule, Eq. 13, with Monte-Carlo sampling to find the optimal control sequence. The algorithm starts by sampling control trajectories, running each trajectory through the dynamics in Eq. 1 to create a corresponding state trajectory, and evaluating each state and control trajectory through the cost function. We provide the option to calculate an importance sampling ratio calculated for the specific sampling distribution like in Eq. 14 or to disable it,
$\mathcal{I}\left(V\right)$ $=\beta\ln\left(\frac{p\left(V\right)}{s\left(V|\theta\right)}\right)$ (16)
where $\beta\in\left\{0,1\right\}$. The option to disable the importance sampling weight comes from experimentation, such as on the AutoRally platform [6], where we have seen better control trajectories without it. Finally, a weighted average of the trajectories is conducted to produce the optimal control trajectory. The update law for $\mathcal{U}^{*}_{t}$, the optimal control at time $t$, ends up looking like
$\mathcal{U}^{*}_{t}$ $=\sum_{m=1}^{M}\frac{\exp\left(-\frac{1}{\lambda}\left(\mathcal{J}\left(V^{m}\right)-\lambda\mathcal{I}\left(V^{m}\right)-\rho\right)\right){\bf v}^{m}_{t}}{\sum_{j=1}^{M}\exp\left(-\frac{1}{\lambda}\left(\mathcal{J}\left(V^{m}\right)-\lambda\mathcal{I}\left(V^{m}\right)-\rho\right)\right)}$ (17)
$\rho$ $=\min_{m\in\left[1,M\right]}\left(\mathcal{J}\left(V^{m}\right)-\lambda\mathcal{I}\left(V^{m}\right)\right)$ (18)
where $V^{m}$ is the $m$-th sampled control trajectory, ${\bf v}^{m}_{t}$ is the control from the $m$-th sampled trajectory at time $t$, and $\rho$ is the minimum sampled cost. When using a Gaussian distribution, the sampled control ${\bf v}^{m}_{t}={\bf u}_{t}+\epsilon^{m}_{t}$ is centered around the previous optimal control ${\bf u}_{t}$ and has noise $\epsilon^{m}_{t}\sim\mathcal{N}\left(0,\Sigma\right)$; other distributions would have different ways to draw samples and alternative $\mathcal{I}\left(V\right)$ equations. We use $\rho$ as in [8] to shift the range of the exponentiated costs to limit numerical stability issues. Sampling in the control space ensures that the trajectories are dynamically feasible and allows us to use non-differentiable dynamics and cost functions. Pseudo-code for the algorithm is shown in Algorithm 1.
Given :  ${\bf F}\left(\cdot,\cdot\right)$, $G\left(\cdot,\cdot\right)$ $\ell\left(\cdot,\cdot\right)$, $\phi\left(\cdot\right)$, $\mathcal{I}\left(\cdot\right)$, $M$, $I$, $T$, $\lambda$, $\Sigma$: System dynamics, system observation, running cost, terminal cost, importance sampling ratio, num. samples, num. iterations, time horizon, temperature, covariance;
Input :  ${\bf x}_{0}$, ${\bf U}$: initial state, mean control sequence;
Output :  $\mathcal{U}$: optimal control sequence
// Begin Cost sampling
1 for $i\leftarrow 1$ to $I$ do
2       for $m\leftarrow 1$ to $M$ do
3             $J^{m}\leftarrow 0$;
4             ${\bf x}\leftarrow{\bf x}_{0}$;
5             for $t\leftarrow 0$ to $T-1$ do
6                   ${\bf v}_{t}\leftarrow{\bf u_{t}}+\epsilon^{m}_{t},\ \epsilon^{m}_{t}\sim\mathcal{N}\left(0,\Sigma\right)$;
7                   ${\bf x}\leftarrow{\bf F}\left({\bf x},{\bf v}_{t}\right)$;
8                   ${\bf y}\leftarrow G\left({\bf x},{\bf v}_{t}\right)$;
9                   $J^{m}\mathrel{+}=\ell\left({\bf y},{\bf v}_{t}\right)-\lambda\mathcal{I}\left({\bf v}_{t}\right)$;
10            $J^{m}\mathrel{+}=\phi\left({\bf y}\right)$
      // Compute trajectory weights
11       $\rho\leftarrow\min\left\{J^{1},J^{2},...,J^{M}\right\}$;
12       $\eta\leftarrow\sum_{m=1}^{M}\exp\left(-\frac{1}{\lambda}\left(J^{m}-\rho\right)\right)$;
13       for $m\leftarrow 1$ to $M$ do
14             $w^{m}\leftarrow\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\left(J^{m}-\rho\right)\right)$;
15
16
      // Control update
17       for $t\leftarrow 0$ to $T-1$ do
18             $\mathcal{U}_{t}\leftarrow{\bf u}_{t}+\sum_{m=1}^{M}w^{m}\epsilon^{m}_{t}$;
19
20
Algorithm 1 MPPI
II-D Library Description
This library is made up of 6 major types of classes:
  • Dynamics
  • Cost Functions
  • Controllers
  • Sampling Distributions
  • Feedback Controllers
  • Plants
The Dynamics and Cost Function classes are self-evident and are classes describing the ${\bf F}$, ${\bf G}$, $\ell$, and $\phi$ functions from Eqs. 1, 19 and 20. The Controller class finds the optimal control sequence ${\bf U}^{*}$ that minimizes the cost in Eq. 20 using algorithms such as MPPI. The Sampling Distributions are used by the Controller class to generate the control samples used for determining the optimal control sequence. The Feedback Controller class determines what feedback controller if any, is used to help push the system back towards the desired trajectory computed by the Controller. This is required for the Tube-MPPI and RMPPI implementations but can even be used with the MPPI controller as it provides feedback in between MPC iterations. Unless otherwise specified, the Feedback Controllers in code examples are instantiated but turned off by default. Finally, Plants are \@iaciMPC MPC wrapper around a given controller and are where the interface methods in and out of the controller are generally defined. For example, a common-use case of MPPI is on a robotics platform running Robot Operating System (ROS) [47]. The Plant is where you would implement your ROS subscribers to information such as state, ROS publishers of the control output, and the necessary methods to convert from ROS messages to MPPI-Generic equivalents. Each class type has their own parameter structures which encapsulate the adjustable parameters of each specific instantiation of the class.
III Performance Implementation
We shall now discuss some of the performance-specific implementation details we make use of in MPPI-Generic. First, we will give a brief introduction to GPU hardware and terminology followed by general GPU performance tricks useful for implementation in this library.
III-A GPU Parallelization Overview
The GPU is a highly parallelizable hardware device that performs operations differently than a CPU. The lowest level of computation in CUDA is a warp, which consists of $32$ threads doing the same operation at every clock step [48]. These warps are grouped together to produce thread blocks. While the threads in a warp are computed together, the individual warps in the thread block are not guaranteed to be at the same place in the code at any given time and sometimes it can actually be more efficient to allow them to differ. The threads in a thread block all have access a form of a local cache called shared memory. Like any memory shared between multiple threads on the CPU, proper mutual exclusion needs to be implemented to avoid race conditions. Threads in a block can be given indices in 3 axes, $x$, $y$, and $z$, which we use to denote different types of parallelization within the library. The conversion from a thread’s 3D index of $(x,y,z)$ to its thread number in the block is given by (z * blockDim.y + y) * blockDim.x + x. Thread blocks can themselves be grouped into grids and are also organized into $x$, $y$, and $z$ axes. This is useful for large parallel operations that cannot fit within a single thread block.The GPU code is compiled into kernels, which can be provided arbitrary grid and block dimensions at runtime.
It is important to briefly understand the hierarchy of memory before discussing how to improve GPU performance. At the highest level, we have global memory which is generally measured in $\mathrm{GB}$s and very slow to access data from. Next, we have an L2 cache in the size of $\mathrm{MB}$s which can speed up access to frequently-used global data. Then we have the L1 cache, shared memory, and CUDA texture caches. The L1 cache and shared memory are actually the same memory on hardware and are generally several $\mathrm{kB}$s in size; they are separated by programmers explicitly using shared memory and the GPU automatically filling the L1 cache. The CUDA texture cache is a fast read-only memory used for CUDA textures which are 2D or 3D representations of data such as a map.
III-B General GPU speedups
When looking into writing more performant code, there are some general tricks that we leveraged throughout our code library. The first is the use of CUDA streams [49]. By default, every call to the GPU blocks the CPU code from moving ahead. CUDA streams allow for the asynchronous scheduling of tasks on the GPU while the CPU performs other work in the meantime. We use CUDA streams throughout in order to schedule memory transfers between the CPU and GPU as well as kernel calls and have different streams for controller optimal control and visualization computations.
The next big tip is minimizing global memory accesses. Global memory reading or writing can be a large bottleneck in computation time and for our library, it can be slower than the actual computations we want to do on the GPU. The first recommendation is to move commonly-accessed data from global memory to shared memory [50]. We also use Curiously Recurring Template Patterns [51] as our choice of polymorphism on the GPU to avoid the need of constructing and reading from a virtual method table which would be stored in global memory.
We utilize vectorized-memory [52] accessing where possible. Looking at the GPU instruction set, CUDA provides instructions to read and write in 32, 64, and 128 bit chunks, making it possible to load up to four $32$-bit floats in a single instruction. Using these concepts, we greatly reduce the number of calls to global memory and consequently increase the speed at which our computations can run.
We also make use of hardware-defined mathematical operators called intrinsics in some places as well to reduce computation time [53]. These instrinsics are approximations of various mathematical operations such as division, $\sin$, etc. that are implemented at the hardware level, requiring many fewer clock cycles to compute. The trade-off is that they are approximations and can return incorrect evaluations depending on the inputs. As such, we limited our use of intrinsics to trigonometric functions like __cosf(), __sinf(), and __sincosf(). We found that other intrinsics such as __fdividef() or __expf() when used throughout the code base cause significantly different optimal control sequence calculations. However, there are plenty of specific locations in the code where more intrinsics can be introduced without negative effects on accuracy at a future point. In addition, we try to make use of float-specific methods when applicable such as expf() to prevent unnecessary conversions to and from double.
Figure 1: Diagram of the execution flow of computeControl(). The blue ellipses indicate variables, the green rectangles are GPU methods, and the orange rectangles are CPU methods. The selection in purple is a single GPU kernel when using the combined kernel and separated out when using split kernels. Most of the code is run on the GPU but we found that some operations such as finding the baseline and calculating the normalizer, $\eta$, run faster on the CPU.
III-C Library-Specific Performance Optimizations
So far, we discussed optimizations that can be done for any CUDA program. However, there are further optimizations to be had in choosing how to parallelize specific components of our library. In Fig. 1, we have the general steps taken every time we want to compute a new optimal control sequence in MPPI. These same steps are also taken in Tube-MPPI and RMPPI though they have to be done for both the nominal and real systems.
One major performance consideration is how to parallelize the Dynamics and Cost Function calculations. We can either run the Dynamics and Cost Function in a combined kernel or run them in separate kernels. We describe each parallelization technique as well as the pros and cons below. First, we introduce a slight modification to the Dynamics and Cost Functions.
III-C1 Intermediate Calculation Passthrough
When creating Cost Functions for a given Dynamics, it might be required to redo calculations already done in the Dynamics. For example, putting a penalty on the location of the wheels of a vehicle inherently requires knowing where the wheels are located. The wheel locations can be calculated given the position and orientation of the center of mass of the vehicle and so are not considered part of the state. Depending on the Dynamics, the wheel locations might also be calculated as part of the state update. To reduce unnecessary recalculations, we provide a way to pass these extra values directly from the Dynamics to the Cost Function. We do this by slightly modifying Eq. 2 to use outputs, ${\bf y}_{t}$ instead of ${\bf x}_{t}$,
${\bf y_{t}}$ $={\bf G}\left({\bf x}_{t},{\bf u}_{t}\right)$ (19)
$\operatorname{\bf J}(Y,U)$ $=\phi({\bf y}_{T})+\sum_{t=0}^{T-1}{\bf\ell}\left({\bf y}_{t},{\bf u}_{t}\right),$ (20)
where ${\bf G}\left({\bf x}_{t},{\bf u}_{t}\right)$ is calculated as part of the Dynamics. For the vast majority of systems, ${\bf y}_{t}$ is the true state, i.e. ${\bf y}_{t}={\bf x}_{t}$, but we have found in some cases that bringing additional calculations from the Dynamics to the Cost Function can be computationally faster than reproducing them.
III-C2 Split Kernel Description
We start by taking the initial state and control samples and run them through the Dynamics kernel. This kernel uses all three axes of thread parallelization for different components. First, the $x$ dimension of the block and the grid are used to indicate which sample we are on as threadIdx.x + blockDim.x * blockIdx.x. As every sample will conduct the exact same computations, using the $x$ axis allows us to ensure that each warp is aligned as long as the $x$ block size is chosen appropriately. Next, the $z$ axis is used to indicate which system is being run; for MPPI, there is only one system but Tube-MPPI and RMPPI use two systems, nominal and real. As dynamics generally have different derivative computations for different states, we use the $y$ dimension, as shown in Lst. 1, to introduce additional parallelization within the dynamics, instead of sequential computation of each state derivative, which can lead to further performance improvements. When our thread block’s $x$ dimension is a multiple of 32, the $y$ threads are separated into different warps and Lst. 1 improves performance. However, when a switch/if statement causes threads in the same warp to follow different computations, this is known as warp divergence, and the GPU runs the warp again to go through all code paths. Depending on the complexity of the branching, this can cause significant slowdowns.
1int tdy = threadIdx.y;
2switch(tdy) {
3 case S_INDEX(X):
4 xdot[tdy] = u[C_INDEX(VEL)]*cos(x[S_INDEX(YAW)]);
5 break;
6 case S_INDEX(Y):
7 xdot[tdy] = u[C_INDEX(VEL)]*sin(x[S_INDEX(YAW)]);
8 break;
9 case S_INDEX(YAW):
10 xdot[tdy] = u[C_INDEX(YAW_DOT)];
11 break;
12}
List of Listings 1 GPU code for the Unicycle Dynamics. This code parallelizes using the thread $y$ dimension to do each state derivative calculation in a different thread
In the Dynamics kernel, we run a for loop over time for each sample in which we get the current control sample, run it through the Dynamic’s step() method, and save out the resulting output to global memory.
Next, we look to the Cost Function ran inside its own kernel. The reason for that is that while the Dynamics must be sequential over time, the cost function does not need to be. To achieve parallelization across time, we move the sample index up to the grid level and use the block’s $x$ axes for time instead. The Cost kernel gets the control and output corresponding to the current time in its computeRunningCost() or terminalCost() methods, adds the cost up across time for each sample, and saves out the resulting overall cost for each sample. A problem that arises is the limited number of timesteps we could optimize over due to the limit of $1024$ threads in a single thread block; we address this by reusing threads to each compute multiple timesteps until we reach the desired time horizon. These choices brings the time to calculate the cost much closer to that of a single timestep instead of having to wait for sequential iterations of the cost if it was paired with the Dynamics kernel.
III-C3 Combined Kernel Description
The Combined Kernel runs the Dynamics and Cost Function methods together in a single kernel. This works by getting the initial state and control samples, applying the Dynamics’ step() to get the next state and output, and running that output through the Cost Functions’ computeCost() to get the cost at that time. This basic interaction is done in a for loop over the time horizon to get the the entire state trajectory as well as the cost of the entire sample. We parallelize this over three axes. First, the $x$ and $z$ dimensions of the block and grid are used to indicate which sample and system we are on as described above in the Split Kernel’s Dynamics section. Finally, the $y$ dimension is used to parallelize within the Dynamics and Cost Functions’ methods.
III-C4 Choosing between the Split and Combined Kernels
There are some trade-offs between the two kernel options that can affect the overall computation time. By combining the Dynamics and Cost Function calculations together, we keep the intermediate outputs in shared memory and do not need to save them out to global memory. However, we are forced to run the Cost Function sequentially in time. Splitting the Dynamics and Cost Function into separate kernels allows them each to use more shared memory for their internal calculations with the requirement of global memory usage to save out the sampled output trajectories. The Combined Kernel uses less global memory but requires more shared memory usage in a single kernel as it must contain both the Dynamics and Cost Functions’ shared memory requirements. As the number of samples grow, the number of reads and writes of outputs to global memory also grows. This can eventually take more time than the savings we get from running the Cost Function in parallel across time, even when using vectorized-memory reads and writes.
In order to address these trade-offs, we implemented both kernel approaches in our library and automatically choose the most appropriate kernel in the Controller constructor using chooseAppropriateKernel(). The automatic kernel selection is done by running both the combined and split kernels multiple times, measuring the computation time of each, and choosing the fastest option. As the combined kernel potentially uses more shared memory than the split kernel, we also check to see if the amount of shared memory used is below the GPU’s shared memory hardware limit; if it is not, we default to the split kernel approach. We also allow the user to overwrite the automatic kernel selection through the use of the setKernelChoice() method.
III-C5 Weight Transform and Update Rule Kernels
Once the costs of each sample trajectory is calculated, we bring these costs back to the CPU in order to find the baseline, $\rho$. The baseline is calculated by finding the minimum cost of all the sample trajectories; it is subtracted out during the exponentiation stage as it has empirically led to better optimization performance. When the number of samples is only in the thousands, we found that the copy to the CPU to do the baseline search is faster than doing so on the GPU. The costs are then exponentiated on the GPU and the normalizer, $\eta$, is calculated back on the CPU before doing the final optimal control calculation.
III-D Performance Recommendations
1#include <mppi/controllers/MPPI/mppi_controller.cuh>
2#include <mppi/cost_functions/cartpole/cartpole_quadratic_cost.cuh>
3#include <mppi/dynamics/cartpole/cartpole_dynamics.cuh>
4#include <mppi/feedback_controllers/DDP/ddp.cuh>
5
6const int NUM_TIMESTEPS = 100;
7const int NUM_ROLLOUTS = 2048;
8using DYN_T = CartpoleDynamics;
9using COST_T = CartpoleQuadraticCost;
10using FB_T = DDPFeedback<DYN_T, NUM_TIMESTEPS>;
11using SAMPLING_T = mppi::sampling_distributions::GaussianDistribution<DYN_T::DYN_PARAMS_T>;
12using CONTROLLER_T = VanillaMPPIController<DYN_T, COST_T, FB_T, NUM_TIMESTEPS, NUM_ROLLOUTS, SAMPLING_T>;
13using CONTROLLER_PARAMS_T = CONTROLLER_T::TEMPLATED_PARAMS;
14
15int main(int argc, char** argv) {
16 float dt = 0.02;
17 std::shared_ptr<DYN_T> dynamics = std::make_shared<DYN_T>(); // set up dynamics
18 std::shared_ptr<COST_T> cost = std::make_shared<COST_T>(); // set up cost
19 // set up feedback controller
20 std::shared_ptr<FB_T> fb_controller = std::make_shared<FB_T>(dynamics.get(), dt);
21 // set up sampling distribution
22 SAMPLING_T::SAMPLING_PARAMS_T sampler_params;
23 std::fill(sampler_params.std_dev, sampler_params.std_dev + DYN_T::CONTROL_DIM, 1.0);
24 std::shared_ptr<SAMPLING_T> sampler = std::make_shared<SAMPLING_T>(sampler_params);
25 // set up MPPI Controller
26 CONTROLLER_PARAMS_T controller_params;
27 controller_params.dt_ = dt;
28 controller_params.lambda_ = 1.0;
29 controller_params.dynamics_rollout_dim_ = dim3(64, DYN_T::STATE_DIM, 1);
30 controller_params.cost_rollout_dim_ = dim3(NUM_TIMESTEPS, 1, 1);
31 std::shared_ptr<CONTROLLER_T> controller = std::make_shared<CONTROLLER_T>(
32 dynamics.get(), cost.get(), fb_controller.get(), sampler.get(), controller_params);
33
34 DYN_T::state_array x = dynamics->getZeroState(); // set up initial state
35 controller->computeControl(x, 1); // calculate control
36 auto control_sequence = controller->getControlSeq();
[…]
38 return 0;
39}
List of Listings 2 Minimal Example to print out optimal control sequence for a cartpole system.
The performance capabilities of the MPPI-Generic library is highly dependent on the choices of block sizes for the various kernels, Dynamics, and Cost Functions. There is no one choice to be made so instead, we provide some general rules of thumbs that we have seen work more often than not.
  • Set the thread block $x$ dimension size of both the Dynamics and Cost kernels to be a multiple of 32. This allows all the threads in a warp to do the same operation. We have seen in some cases that lowering the thread block $x$ dimension to 16 can provide some improvements but this has not been true for the majority of dynamics and cost functions.
  • Set the thread block $y$ dimension for Dynamics equal to the state dimension and Cost Functions to 1. This depends on whether the Dynamics/Cost Function are utilizing the multi-threading capability available to them but for most of the pre-defined Dynamics, they are using the $y$ axis of the thread block to do dynamics updates. Meanwhile, most Cost Functions are not taking advantage of the parallelization as they return a single value, the cost.
  • Keep Dynamics and Cost Function kernels’ thread block sizes low. The limit to the size of a thread block is currently $1024$ so it might be tempting to fill that. However, there are various synchronization points in these kernels to ensure that data has been loaded before being used. As the thread block size increases, more time is spent waiting at these synchronization points in the Dynamics and Cost Function kernels.
  • When developing your own Dynamics, it can be pertinent to make the output dimensions divisible by 2 or 4. This can done by adding extraneous values to the enum. This allows the code to use more efficient memory-loading GPU instructions discussed in Section III-B and reduce the number of memory calls by a factor of at least 2. One might be tempted to also ensure that the control dimension is divisible by 4. However, the control dimension is also used to generate samples so while there may be improvements to the efficiency of reading memory, it will also slow down generating samples.
  • Minimize the occurrence of warp divergence in your GPU code when possible. The most common way these occur is by using if statements based on values that might be different for individual threads in a warp. In some cases, converting the desired bool expression into a float will be faster while still providing the desired branching.
IV API Structure
We describe the library API and its usage in three stages. At the beginner level, the desire is to use a provided Dynamics, Cost Function, and Controller to control a system. This requires the least amount of code writing on the part of the user as most of the code is already provided. The library user would only need to write a specific Plant class to properly interface with whatever system they are wanting to control and the executable which sets up and runs the controller itself. The intermediate level is where the user might want to implement a dynamics model or cost function that does not exist in the base library. Finally, at the advanced level, we show how to implement a custom Feedback Controller, Controller, or Sampling Distribution.
IV-A Beginner
1#pragma once
2#include <mppi/core/base_plant.hpp>
3#include <mppi/dynamics/cartpole/cartpole.cuh>
4
5template <class CONTROLLER_T> class SimpleCartpolePlant : public class BasePlant<CONTROLLER_T> {
6public:
7 using control_array = typename CartpoleDynamics::control_array;
8 using state_array = typename CartpoleDynamics::state_array;
9 using output_array = typename CartpoleDynamics::output_array;
10
11 SimpleCartpolePlant(std::shared_ptr<CONTROLLER_T> controller, int hz, int optimization_stride)
12 : BasePlant<CONTROLLER_T>(controller, hz, optimization_stride) {
13 system_dynamics_ = std::make_shared<CartpoleDynamics>();
14 }
15
16 void pubControl(const control_array& u) {
17 state_array state_derivative;
18 output_array dynamics_output;
19 state_array prev_state = current_state_;
20 float t = this->state_time_;
21 float dt = this->controller_->getDt();
22 system_dynamics_->step(prev_state, current_state_, state_derivative, u, dynamics_output, t, dt);
23 current_time_ += dt;
24 }
25
26 void pubNominalState(const state_array& s) {}
27 void pubFreeEnergyStatistics(MPPIFreeEnergyStatistics& fe_stats) {}
28 int checkStatus() { return 0; }
29 double getCurrentTime() { return current_time_; }
30 double getPoseTime() { return this->state_time_; }
31
32 state_array current_state_;
33protected:
34 std::shared_ptr<CartpoleDynamics> system_dynamics_;
35 double current_time_ = 0.0;
36}
List of Listings 3 Basic Plant implementation that interacts with a virtual Cartpole dynamics system stored within the Plant.
We start by showing an example of just using a single iteration of MPPI to produce an optimal control sequence in Lst. 2. At the beginning of the example (Lines 8, 9, 10, 11, 12 and 13), we create aliases such DYN_T for CartpoleDynamics to keep the code fairly succinct. The Feedback Controller is not used for this example but it is required to exist. The Sampling Distribution and Controller have parameters that need to be set; their use spans from algorithmic to performance – Lines 22, 23, 26, 27, 28, 29 and 30 respectively. The std_dev, dt_, and lambda_ are parameters affecting the MPPI update rule whereas the dynamics_rollout_dim_ and cost_rollout_dim_ parameters adjust how fast the MPPI algorithm is computed using recommendations from Section III-D. Once the components are initialized, we create an instance of the MPPI Controller, passing the Dynamics, Cost Function, Feedback Controller, Sampling Distribution, and controller parameters to the constructor. On Line 34, we create an initial zero state for the dynamics using getZeroState() and compute an optimal control sequence with computeControl() on Line 35. We return the control sequence as an Eigen::Matrixf with $n_{u}$ rows and $T$ columns using getControlSeq() to print out.
When using MPPI in \@iaciMPC MPC fashion, we need to use a Plant wrapper around our controller. The Plant houses methods to obtain new data such as state, calculate the optimal control sequence at a given rate using the latest information available, and provide the latest control to the external or ground truth system while providing the necessary tooling to ensure there are no race conditions. As this class provides the interaction between the algorithm and the actual system, it is a component that has to be modified for every use case. For Lst. 3, we implement a plant (SimpleCartpolePlant) inheriting from BasePlant that contains the ground truth system completely internal to the class. Specifically, our plant runs the external dynamics inside pubControl() in order to produce a new state. We then call updateState() at a different fixed rate from the controller re-planning rate to show that the capability of the code base. SimpleCartpolePlant instantiates a CartpoleDynamics object in its constructor, overwrites the required virtual methods from BasePlant, and sets up the dynamics update to occur within pubControl(). Looking at the constructor on Line 11, we pass a shared pointer to a Controller, an integer representing the controller replanning rate, and the minimum optimization stride, before creating our stand-in system dynamics. In a new MPC iteration, we shift the start of the mean control sequence we sample around to the maximum between the number of timesteps since the last optimization and the minimum optimization stride. For use in MPC, we recommend setting the minimum optimization stride to $1$. pubControl() on Line 16 is where we send the control to the system. In this case, we create necessary extra variables to pass the current state ${\bf x}_{t}$ and control ${\bf u}_{t}$ as prev_state and u respectively to the Dynamics’ step() method to get the next state, ${\bf x}_{t+1}$, in the variable current_state_. We also update the current time on Line 23 to show the system has moved forward in time. Looking at this class, an issue arises as the Controller it is templated upon might not use CartpoleDynamics as its Dynamics class. This is easily remedied by replacing any reference to CarpoleDynamics with CONTROLLER_T::TEMPLATED_DYNAMICS to make this Plant work with the Dynamics used by the instantiated Controller.
Now that we have written our specialized Plant class, we can make some modifications to Lst. 2 to use the controller in \@iaciMPC MPC fashion. For this example, we would run a simple for loop that calls the Plant’s runControlIteration() and updateState() methods to simulate a receiving a new state from the system and then calculating a new optimal control sequence from it, replacing Line 35. The updateState() method calls pubControl() internally so the system state and the current time would update at each for loop iteration. For real-time scenarios, the runControlLoop() Plant method can be launched in a separate thread and calls runControlIteration() internally at the specified re-planning rate.
IV-B Intermediate
In the previous section, much of the underlying structure of the library was glossed over. As we get to implementing our own Dynamics or Cost Functions however, there are some basic principles to go over. This library is running code on two different devices, the CPU and the GPU. Some classes, such as Plants and Controllers, do not have methods that need to run on both whereas other classes like the Dynamics and Cost Functions do. The GPU can do many computations in parallel but in order to properly utilize it, that can require different code than what runs on the CPU. As such, when looking at implementing a new Dynamics or Cost Function, there are some methods that have to be implemented as two very similar functions, once for the CPU and once for the GPU.
IV-B1 Custom Dynamics
The first thing that needs to be implemented for a new Dynamics class is a new parameter structure. We implemented a dictionary-like structure in the form of enum to define the state, control, and output vectors and these dictionaries are stored in the parameter. For example, in Lst. 4, we show a basic parameter structure implemented for a unicycle model.
6struct UnicycleParams : public DynamicsParams {
7 enum class StateIndex : int {
8 X = 0,
9 Y,
10 YAW,
11 NUM_STATES
12 };
13
14 enum class ControlIndex : int {
15 VEL = 0,
16 YAW_DOT,
17 NUM_CONTROLS
18 };
19
20 enum class OutputIndex : int {
21 X = 0,
22 Y,
23 YAW,
24 NUM_OUTPUTS
25 };
26};
List of Listings 4 Simple parameter structure implementation for a unicycle model
By implementing these enum, we can use these state names later on in the Dynamics and Cost Function to make it clear what state we are referring to. The ending values of NUM_STATES, NUM_CONTROLS, and NUM_OUTPUTS are also used to determine the size of each of the resulting dimensions for creating statically-sized Eigen::Matrixf data types such as state_array and control_array.
28class Unicycle : public Dynamics<Unicycle, UnicycleParams> {
29public:
30 using PARENT_CLASS = Dynamics<Unicycle, UnicycleParams>;
31
32 std::string getDynamicsModelName() const override {
33 return ”Unicycle”;
34 }
35
36 Unicycle(cudaStream_t stream = nullptr)
37 : PARENT_CLASS(stream) {}
38
39 void computeStateDeriv(
40 const Eigen::Ref<const state_array>& x,
41 const Eigen::Ref<const control_array>& u,
42 Eigen::Ref<state_array> x_dot);
43
44 __device__ inline void computeStateDeriv(
45 float* x, float* u,
46 float* x_dot, float* theta_s);
47
48 state_array stateFromMap(const std::map<std::string, float>& map) override;
49};
List of Listings 5 All the methods that need to overwritten in a custom Dynamics class
Next, we implement the necessary overwritable methods. These methods are shown in Lst. 5 and start with getDynamcisModelName() which returns the name of the Dynamics model. Next are the CPU and GPU versions of computeStateDeriv() on Lines 39 and 44 respectively. The CPU and GPU versions are differentiated firstly by the __device__ keyword at the front of the GPU code to designate that method only runs on the GPU. Next, we also only use Eigen::Matrixf data-types on the CPU and raw float pointers on the GPU. The computeStateDeriv() both would implement the following dynamics,
$\dot{x}$ $=u_{0}*cos\left(\psi\right)$ (21)
$\dot{y}$ $=u_{0}*sin\left(\psi\right)$ (22)
$\dot{\psi}$ $=u_{1},$ (23)
as shown in Lst. 6. Note the use of the enum we created earlier with the S_INDEX() and C_INDEX() macros. This allows us to not need to know the order of the states or controls in the underlying data type and more importantly, these same enum can also be used in Cost classes to ensure that compatibility with multiple dynamics as long as the dynamics define the necessary enum values.
5xdot[S_INDEX(X)] = u[C_INDEX(VEL)] * cos(x[S_INDEX(YAW)]);
6xdot[S_INDEX(Y)] = u[C_INDEX(VEL)] * sin(x[S_INDEX(YAW)]);
7xdot[S_INDEX(YAW)] = u[C_INDEX(YAW_DOT)];
List of Listings 6 computeStateDeriv() implementation for a Unicycle dynamics
An important note is that there are separate instances of the class allocated on the CPU and GPU for dynamics and cost functions. In order to update the GPU version, you must overwrite paramsToDevice() to copy from the CPU side class to the GPU side class. Parameter structures have a default implementation for this, but if helper classes are added, their corresponding copy methods will need to be called.
Finally, the last method to be overwritten is stateFromMap() This method is used to translate std::map of state names and values into the Dynamics’ corresponding state vector. This method ends up being useful for the Plant, especially when different parts of the state can come in at different rates.
The methods listed above are just the beginning of the Dynamics API customization to get started. More advanced customization includes changing the default choice of integration scheme from Euler integration to more accurate integration schemes such as Runge-Kutta or implicit integration if needed, adjusting what is considered the zero state and control for the Dynamics, and adjusting how to linearly interpolate between states.
IV-B2 Custom Cost Function
6struct UnicycleCostParams : public CostParams<Unicycle::CONTROL_DIM>
7{ float width = 1.0; float coeff = 10.0; };
8
9class UnicycleCost : public Cost<UnicycleCost, UnicycleCostParams, Unicycle::DYN_PARAMS_T>
10{
11public:
12 using PARENT_CLASS = Cost<UnicycleCost, UnicycleCostParams, Unicycle::DYN_PARAMS_T>;
13 using DYN_P = PARENT_CLASS::TEMPLATED_DYN_PARAMS;
14
15 UnicycleCost(cudaStream_t stream = nullptr);
16
17 std::string getCostFunctionName()
18 { return ”Unicycle Cost”; }
19
20 float computeStateCost(
21 const Eigen::Ref<const output_array> y,
22 int t, int* crash_status);
23
24 float terminalCost(
25 const Eigen::Ref<const output_array> y);
26
27 __device__ float computeStateCost(float* y,
28 int t, float* theta_c, int* crash_status);
29
30 __device__ float terminalCost(float* y,
31 float* theta_c);
32};
List of Listings 7 Basic Cost Function Parameter Structure and Class Implementation for the Unicycle Dynamics
If we make a Dynamics class, we also generally need to make a corresponding Cost. While a default quadratic cost implementation that allows for use with any dynamics exists, we implement a quick Cost class to show what methods to overwrite. For this example, we choose to think of our unicycle on a road of some width pointing in the $x$ direction. We want to keep the unicycle on the road so we penalize the absolute $y$ position linearly up to the road width and penalize it quadratically if it goes outside the road.
To do this, we first construct a parameter structure for the cost class. It inherits from the base CostParams class and needs to know the CONTROL_DIM and have a variable for the width of the road and coefficient for the cost. From there, we implement a basic Cost class that has to overwrite computeStateCost() and terminalCost() on both the CPU and GPU sides. It also creates a default constructor and name method getCostFunctionName(). This basic implementation is shown in Lst. 7. Notice that we created an alias for our dynamics parameter’s struct on Line 13. This is to allow us to use the enum defined there inside our cost function as well. Other things to note are that again, our cost is based on the output rather than the state in computeStateCost(). For our simple Unicycle Dynamics, the output is the same as the state.
9 float cost = 0;
10 float y_abs = abs(y[O_IND_CLASS(DYN_P, Y)]);
11 if (y_abs < this->params_.width) {
12 cost = y_abs;
13 } else { // Quadratic cost outside the road width
14 cost = y_abs * y_abs;
15 }
16 return this->params_.coeff * cost;
List of Listings 8 Basic State Cost Implementation for the Unicycle Dynamics
In Lst. 8, the cost function described previously is implemented. We can use enum macros such as O_IND_CLASS() when we are outside of the Dynamics class to still find the appropriate output value. This code can be implemented in both the CPU and GPU computeStateCost() methods and we choose terminalCost() to return 00.
These new Dynamics and Cost are easily incorporated into our previous controller examples and that should be enough to get most people started on using this library. There are more options in the Cost and Dynamics classes to improve GPU performance but those are left to Section III.
IV-C Advanced
In this final subsection, we show off how to customize Controllers, Feedback Controllers, and Sampling Distributions. Customizing these allows users to create new sampling-based controllers or change the sampling distributions used in specific scenarios.
IV-C1 Feedback Controllers
Figure 2: Feedback Controller API Diagram. We have a GPUState-based parameter structure that contains things like the feedback gains of iLQR. The GPU Feedback Controller class exists both on the GPU (shown in green) and CPU (shown in orange) and contains the GPUState as well as a method $k(x,x^{*})$ to calculate the feedback control on the GPU. The CPU Feedback Controller class (shown in yellow) is a wrapper around the GPU Feedback class that has a CPU method to calculate $k(x,x^{*})$ as well as a method, computeFeedback(), to recompute the feedback gains.
Feedback Controllers are useful even when using MPPI and become necessary for Tube-MPPI and RMPPI. The only Feedback Controller currently implemented is iLQR but we show how to construct a simple PID controller for each state/control combination.
7template <class DYN_T>
8class PIDState : public GPUState
9{
10 using STATE_DIM = DYN_T::STATE_DIM;
11 using CONTROL_DIM = DYN_T::CONTROL_DIM;
12 float p[STATE_DIM * CONTROL_DIM] = {0.0};
13 float i[STATE_DIM * CONTROL_DIM] = {0.0};
14 float d[STATE_DIM * CONTROL_DIM] = {0.0};
15 float dt;
16};
List of Listings 9 PID Controller Parameter Structure containing the $p$, $i$, and $d$ feedback matrices as well as the $\Delta t$ used for integral and derivative calculations.
Before the example Feedback Controller, we need to go over the code structure for Feedback Controllers as this is very different from the previous examples of Dynamics and Cost Functions. This is due to the fact that the feedback controller needs to be usable on the GPU but might have different memory requirements on the CPU side. Using iLQR as an example in Fig. 2, calculating the feedback on the GPU would just require the feedback gains to be sent but calculating the feedback gains requires access to the Dynamics and their Jacobians as well as a Cost Function and its Jacobian. To handle this discrepancy in computational workloads, each Feedback Controller is made up of two parts: a GPU feedback class which contains the bare necessities to calculate the feedback, and a CPU wrapper which can do things like recomputing the gains to a new desired trajectory.
Now let us show how to utilize the Feedback Controller API to create a new PID controller. We start by constructing the parameter structure for the GPU Feedback class, shown in Lst. 9. Note that it is templated off of the Dynamics so that it can create the appropriately-sized arrays for each gain. Since this is a generic PID controller, we use the full feedback matrix representation for each type of gain. Next, we need to implement the GPU Feedback Controller for the PID in Lst. 10. The only method that is required to be overwritten is the k() feedback method on Line 23 but as PIDs require some history to calculate the $i$ and $d$ portions, we request extra memory by setting the SHARED_MEM_REQUEST_BLK_BYTES variable on Line 21. This variable allows for extra shared memory per trajectory sample which is necessary for keeping track of history on the GPU.
18template <class DYN_T>
19class gpuPID : public GPUFeedbackController<gpuPID<DYN_T>, DYN_T, PIDState<DYN_T>> {
20public:
21 static const int SHARED_MEM_REQUEST_BLK_BYTES = DYN_T::STATE_DIM * 2;
22
23 __device__ void k(const float* x_act,
24 const float* x_goal, const int t,
25 float* theta, float* u_fb) {
26 int tid = threadIdx.x + blockDim.x * threadIdx.z;
27 float* e_int = theta[(tid*2+0) * DYN_T::STATE_DIM];
28 float* pre_e = theta[(tid*2+1) * DYN_T::STATE_DIM];
29 float e_der[DYN_T::STATE_DIM];
30 float curr_e[DYN_T::STATE_DIM];
31 for (int i = 0; i < DYN_T::STATE_DIM; i++) {
32 curr_e[i] = (x_act[i] - x_goal[i]);
33 e_der[i] = (curr_e[i] - pre_e[i]) / this->state_.dt;
34 e_int[i] += curr_e[i] * this->state_.dt;
35 pre_e[i] = curr_e[i];
36 }
37
38 // start calculating control
39 for (int i = 0; i < DYN_T::CONTROL_DIM; i++) {
40 for (int j = 0; j < DYN_T::STATE_DIM; j++) {
41 int fb_idx = i + j * DYN_T::CONTROL_DIM;
42 u_fb[i] = this->state_.p[fb_idx] * curr_e[j];
43 u_fb[i] += this->state_.i[fb_idx] * e_int[j];
44 u_fb[i] += this->state_.d[fb_idx] * e_der[j];
45 }
46 }
47}
48};
List of Listings 10 PID GPU Implementation showing how to implement $k(x,x^{*})$ and request shared memory for the integral and derivative of states.
51template <class DYN_T, int NUM_TIMESTEPS>
52class PIDFeedback : public FeedbackController<gpuPID<DYN_T>, PIDParams, NUM_TIMESTEPS> {
53public:
54 using PARENT_CLASS = FeedbackController<gpuPID<DYN_T>, PIDParams, NUM_TIMESTEPS>;
55 using control_array = typename PARENT_CLASS::control_array;
56 using state_array = typename PARENT_CLASS::state_array;
57 using state_trajectory = typename PARENT_CLASS::state_trajectory;
58 using control_trajectory = typename PARENT_CLASS::control_trajectory;
59 using INTERNAL_STATE_T = typename PARENT_CLASS::TEMPLATED_FEEDBACK_STATE;
60 using feedback_gain_matrix = typename DYN_T::feedback_matrix;
61
62
63 PIDFeedback(float dt = 0.01, int num_timesteps = NUM_TIMESTEPS, cudaStream_t stream = 0)
64 : PARENT_CLASS(dt, num_timesteps, stream) {
65 this->gpu_controller_->getFeedbackStatePointer()->dt = dt;
66 }
67
68 control_array k_(const Eigen::Ref<const state_array>& x_act, const Eigen::Ref<const state_array>& x_goal,
69 int t, INTERNAL_STATE_T& fb_state) {
70 Eigen::Map<feedback_gain_matrix> P_gain(&fb_state.p);
71 Eigen::Map<feedback_gain_matrix> I_gain(&fb_state.i);
72 Eigen::Map<feedback_gain_matrix> D_gain(&fb_state.d);
73 state_array curr_e = x_act - x_goal;
74 state_array der_e = (curr_e - prev_e) / this->dt_;
75 e_int_ += (curr_e) * this->dt_;
76 prev_e_ = curr_e;
77 control_array output = P_gain * curr_e + I_gain * e_int_ + D_gain * der_e;
78 return output;
79 }
80
81 void computeFeedback(const Eigen::Ref<const state_array>& init_state,
82 const Eigen::Ref<const state_trajectory>& goal_traj,
83 const Eigen::Ref<const control_trajectory>& control_traj) {}
84
85protected:
86 state_array e_int_ = state_array::Zero();
87 state_array prev_e_ = state_array::Zero();
88};
List of Listings 11 PID CPU Implementation. It shows the necessary type aliases and how to compute the feedback control in k_().
After writing the GPU feedback class, we now just have to write the CPU feedback class, shown in Lst. 11. We have two methods to overwrite here in computeFeedback() and k_(). computeFeedback() is used to calculate new feedback gains given a new trajectory. For our simple PID class, we stick to constant PID gains so this can be left empty. The k_() method on Line 68 is just the CPU version of the feedback calculation. For the CPU version of the PID, we can now use member variables to hold the history required to calculate the $i$ and $d$ portions.
IV-C2 Controller
13template <int STATE_DIM, int CONTROL_DIM, int MAX_T>
14struct CEMParams : public ControllerParams<STATE_DIM, CONTROL_DIM, MAX_T> {
15 float top_k_percentage = 0.10f;
16};
List of Listings 12 CEM Controller Parameter Structure containing the percentage of elite samples.
48void computeControl(const Eigen::Ref<const state_array>& state, int optimization_stride = 1) override {
49 // Send the initial condition to the device
50 HANDLE_ERROR(cudaMemcpyAsync(this->initial_state_d_, state.data(), DYN_T::STATE_DIM * sizeof(float),
51 cudaMemcpyHostToDevice, this->stream_));
52 for (int opt_iter = 0; opt_iter < this->getNumIters(); opt_iter++) {
53 this->copyNominalControlToDevice(false); // Send the optimal control sequence to the device
54 this->sampler_->generateSamples(optimization_stride, opt_iter, this->gen_, false); // Generate noise data
55
56 // Calculate state trajectories and costs from sampled control trajectories
57 mppi::kernels::launchSplitRolloutKernel<DYN_T, COST_T, SAMPLING_T>(
58 this->model_->model_d_, this->cost_->cost_d_, this->sampler_->sampling_d_, this->getDt(),
59 this->getNumTimesteps(), NUM_SAMPLES, this->getLambda(), this->getAlpha(), this->initial_state_d_,
60 this->output_d_, this->trajectory_costs_d_, this->params_.dynamics_rollout_dim_,
61 this->params_.cost_rollout_dim_, this->stream_, false);
62 // Copy the costs back to the host
63 HANDLE_ERROR(cudaMemcpyAsync(this->trajectory_costs_.data(), this->trajectory_costs_d_,
64 NUM_SAMPLES * sizeof(float), cudaMemcpyDeviceToHost, this->stream_));
65 HANDLE_ERROR(cudaStreamSynchronize(this->stream_));
66
67 // Setup vector to hold top k weight indices
68 int top_k_to_keep = NUM_SAMPLES * this->params_.top_k_percentage;
69 calculateEliteSet(this->trajectory_costs_, NUM_SAMPLES, top_k_to_keep, top_k_indices_);
70 // keep weights of the elite set
71 float min_elite_value = this->trajectory_costs_[top_k_indices_.back()];
72 std::replace_if(
73 this->trajectory_costs_.data(), this->trajectory_costs_.data() + NUM_SAMPLES,
74 [min_elite_value](float cost) { return cost < min_elite_value; }, 0.0f);
75
76 // Copy weights back to device
77 HANDLE_ERROR(cudaMemcpyAsync(this->trajectory_costs_d_, this->trajectory_costs_.data(),
78 NUM_SAMPLES * sizeof(float), cudaMemcpyHostToDevice, this->stream_));
79 // Compute the normalizer
80 this->setNormalizer(mppi_common::computeNormalizer(this->trajectory_costs_.data(), NUM_SAMPLES));
81 // Calculate the new mean
82 this->sampler_->updateDistributionParamsFromDevice(this->trajectory_costs_d_, this->getNormalizerCost(),
83 0, false);
84 // Transfer the new control back to the host and synchronize stream
85 this->sampler_->setHostOptimalControlSequence(this->control_.data(), 0, true);
86 }
87 // Calculate optimal state and output trajectory from the current state and optimal control
88 this->computeOutputTrajectoryHelper(this->output_, this->state_, state, this->control_);
89}
List of Listings 13 Basic CEM computeControl() implementation. This method copies the initial state to the GPU, creates the control samples, calculates the state trajectories and costs of each sample, creates the elite set, updates the mean of the sampling distribution and copies that mean back as the optimal control sequence.
We walk through how to create a new sampling-based Controller using CEM [5] as our desired algorithm. Succinctly put, the CEM method samples control trajectories from a Gaussian Distribution, and uses the best $k$ samples, known as the elite set, to calculate new parameters for the Guassian distribution. For this example, we simplify it to have constant variance and the elite set is used to update the mean of the distribution. The major CEM parameter is the size of the elite set, so we create a new parameter struct in Lst. 12 capturing this as a percentage of the total number of samples. From there, the methods that must be overwritten from the base Controller API are getControllerName(), computeControl(), and calculateSampledStateTrajectories(). Like in the other custom classes, getControllerName() returns the name of the new controller, i.e. CEM. Next, computeControl() is the main method of the Controller class. It takes the new initial state and calculates the new optimal control sequence from that starting position; an example is shown in Lst. 13. The basic steps are to move the shifted optimal control sequence and initial state to the GPU, generate control samples, run the samples through the Dynamics and Cost function, calculate the weights of each sample, use these weights to update the parameters of the sampling distribution, get the new optimal control sequence, and calculate the corresponding optimal state trajectory. In Lines 68, 69, 70, 71, 72, 73 and 74, we add an additional method, calcualteEliteSet(), to find the elite set and zero out the weights of every other sample specifically for CEM. The final method to overwrite, calculateSampledStateTrajectories(), is a method used to return a subset of sampled trajectories from the latest optimization round from the GPU to the CPU. Users can set a percentage of the sampled trajectories they would like returned with setPercentageSampledControlTrajectoriesHelper() and this method will generate state trajectories for those samples so that they can then be used for visualization in programs such as RViz. For this simple example, we have no visualization system to plug this data into so we leave this method empty for our CEM implementation.
IV-C3 Sampling Distributions
The Sampling Distribution class is where the choice of how to sample the control distributions is conducted. Different sampling distributions can have significant impact to the controller performance and is still a large area of research being explored. The current sampling distributions implemented are Gaussian, Colored Noise [20], Normal log-Normal (NLN) noise [21], and Smooth-MPPI [42], but even in those implementations, there are further options and tweaks that we found to improve performance in the past. Some of these include having a percentage of samples sampled from a zero-mean distribution rather than the previous optimal control sequence, using the mean with no noise as a sample, allowing for time-varying standard deviations, and the ability to disable the importance sampling weight.
The Sampling Distribution API leaves enough flexibility to allow for multi-hypothesis distributions such as Gaussian Mixture Model (GMM) distributions or Stein-Variational distributions [43, 54]. In addition, the readControlSample() method used to get the control sample for a particular sample, time, and system takes in the current output which allows for feedback-based sampling such as done in normalizing flow approaches [55, 56]. The essential methods to focus on when implementing a new Sampling Distribution are generateSamples() and readControlSample().
V Benchmarks
In order to see the improvements our library can provide, we decided to compare against three other implementations of MPPI publicly available. The first comparison is with the MPPI implementation in AutoRally [57]. This implementation was the starting point of our new library, MPPI-Generic, and so we want to compare to see how well we can perform to our predecessor. The Autorally implementation is written in C++/CUDA, is compatible with ROS, features multiple dynamics models including linear basis functions, simple kinematics, and Neural Network (NN)-based models focused on the Autorally hardware platform. There is only one Cost Function available but it makes use of CUDA textures querying into an obstacle map. Additionally, it has been shown to run in real-time on hardware to great success [6, 4, 8]. However, the Autorally implementation is written for use on the Autorally platform and has no general Cost Function, Dynamics, or Sampling Distribution APIs to extend. In order to use it for different problems such as flying a quadrotor, the MPPI implementation would need significant modification.
The next implementation we compare against is ROS2’s MPPI. As of ROS Iron, there is a CPU implementation of MPPI in the ROS navigation stack [58]. This CPU implementation is written in C++ and looks to make heavy use of AVX or vectorized instructions to improve performance. There is a small selection of dynamics models (Differential Drive, Ackermann, and Omni-directional) and cost functions that are focused around wheeled robots navigating through obstacle-laden environments. This implementation will only become more widespread as ROS2 adoption continues to grow over the coming years, making it an essential benchmark. Unfortunately, it does have some drawbacks as it is not possible to add new dynamics or cost functions without rewriting the base code itself, has no implementation of Tube-MPPI or RMPPI, and is only available in ROS2 Iron or newer. This means that it might not be usable on existing hardware platforms that are unable to upgrade their systems.
The last implementation of MPPI we compare against is in TorchRL [59]. TorchRL is an open-source Reinforcement Learning (RL) Python library written by Meta AI, the developers of PyTorch itself. As such, it is widely trusted and available to researchers who are already familiar with PyTorch and Python. The TorchRL implementation works on both CPUs and GPUs and allows for custom dynamics and cost functions through the extension of base Environment class [60]. However, while it does have GPU support, it is limited to the functionality of PyTorch meaning that there is no option to use CUDA textures to improve map queries or any direct control of shared memory usage on the GPU. In addition, being written in Python makes it fairly legible and easy to extend but can come at the cost of performance when compared to C++ implementations.
In order to compare our library against these three implementations, we recreated the same dynamics and cost function for each version of MPPI. As ROS2’s implementation would be the hardest to modify, we chose the Differential Drive dynamics model and some of the cost function components that already exist there as the baseline. We used the goal position quadratic cost, goal angle quadratic cost, and the costmap-based obstacle cost components so that we could maintain a fairly simple cost function that allows us to show the capabilities of our library. We implemented these dynamics and cost functions in both CUDA and Python. The CUDA implementations were extensions of our base Dynamics and Cost Function APIs. We decided to use the same code in the Autorally implementation as well which required some minor rewriting to account for different method names and state dimensions. The Python implementation was an extension of the TorchRL base Environment class, and used PyTorch’s JIT compiler to speed up performance when used in the TorchRL implementation. We used the same parameters for sampling, dynamics, cost function tuning, and MPPI hyperparameters across all implementations, summarized in Table I.
TABLE I: Algorithm Parameters
Parameter Value
dt 0.02 s
wheel radius 1.0 $\mathrm{m}$
wheel length 1.0 $\mathrm{m}$
max velocity 0.5 $\mathrm{m}\mathrm{/}\mathrm{s}$
min velocity -0.35 $\mathrm{m}\mathrm{/}\mathrm{s}$
min rotation -0.5 $\mathrm{rad}\mathrm{/}\mathrm{s}$
max rotation 0.5 $\mathrm{rad}\mathrm{/}\mathrm{s}$
MPPI Parameters
$\lambda$ 1.0
control standard deviation 0.2
MPC Horizon 100 timesteps
Cost Parameters
Dist. to goal coefficient 5
Angular Dist. to goal coeff 5
Obstacle Cost 20
Map width 11 $\mathrm{m}$
Map Height 11 $\mathrm{m}$
Map Resolution 0.1 $\mathrm{m}\mathrm{/}\mathrm{\text{cell}}$
For both Autorally and MPPI-Generic, there are further performance-enhancing options available such as block size choice. We ended up using the same block sizes for both Autorally and MPPI-Generic across all tests, shown in Table II.
TABLE II: GPU Performance Choices
Parameter Value
Dynamics thread block x dim. 64
Dynamics thread block y dim. 4
Cost thread block x dim. 64
Cost thread block y dim. 1
As a result, the optimization times shown are not going to be the fastest possible performance that can be achieved on any given GPU but these tests should still serve as a useful benchmark to understand the average performance that can be achieved. Our test was timing how long each implementation of MPPI would take to return an optimal control sequence when provided an initial state, ${\bf x}_{0}$. We ran each of the Autorally, MPPI-Generic, and ROS2 implementations $10,000$ times to produce optimal trajectories with 128, 256, 512, 1024, 2048, 4096, 6144, 8192, and 16,384 samples; the TorchRL implementation was only run $1000$ times due to it being too slow to compute, even when using the GPU. The comparisons were run across a variety of hardware including a Jetson Nano to see what bottlenecks each implementation might have. The Jetson Nano was unfortunately only able to run the MPPI-Generic and Autorally MPPI implementations as the last supported PyTorch version and the lastest TorchRL libraries were incompatible, and the ROS2 implementation was unable to compile. GPUs tested ranged from a NVIDIA GTX 1050 Ti to a NVIDIA RTX 4090. Most tests were performed on an Intel 13900K which is one of the fastest available CPUs at the time of this writing in order to prevent the CPU being the bottleneck for the mostly GPU-based comparison; however, we also ran tests on an AMD Ryzen 5 5600x to see the difference in performance on a lower-end CPU. The MPPI optimization times across all hardware can be seen in Table III. The code used to do these comparisons is available at https://github.com/ACDSLab/MPPI_Paper_Example_Code.
V-A Results
While this results section focuses on the computational speed of our algorithm for fairly simple dynamics and cost functions, readers may be curious to know how this library performs in real-world applications. Adding those results to this paper would end up requiring too much space for this publication; instead, we point readers to other papers where this code base has been used in more complex scenarios [20, 24, 11, 33, 34].
Figure 3: Optimization times for all MPPI implementations on a hardware system with a RTX 4090 and an Intel 13900K over a variety of number of samples. The ROS2 CPU implementation grows linearly as the number of samples increase while GPU implementations grow more slowly.
Going over all of the collected data would take too much room for this paper so we shall instead try to pull out interesting highlights to discuss. Full results can be seen in Table III. First, we look at the results on the most powerful system tested, using an Intel i9-13900K and an NVIDIA RTX 4090 in Fig. 3. As the number of samples increase, the CPU-bound ROS2 method increases in optimization times in a linear fashion. Every other method uses the GPU and we see little reason to use small sample sizes as they have the same computation time till we reach around 1024 samples. As we hit $16,384$ samples, the AutoRally implementation starts to have lower optimization times than MPPI-Generic. We will see this trend continue in Fig. 4.
Figure 4: Optimization times for all MPPI implementations on a hardware system with a GTX 1050 Ti and an Intel 13900K over a variety of number of samples. MPPI-Generic and AutoRally on this older hardware eventually start to scale linearly with the number of samples but does so at a much lower rate with our library compared to ROS2 or TorchRL.
When looking at older and lower-end NVIDIA hardware such as the GTX 1050 Ti, our library still performs well compared to other implementations as seen in Fig. 4. Only when the number of samples is at 128 does the ROS2 implementation on an Intel 13900k match the performance of the AutoRally implementation on this older GPU. MPPI-Generic is still more performant at these lower number of samples and eventually it scales linearly as we get to thousands of samples. The TorchRL implementation also finally starts to show some GPU bottle-necking as we start to see optimization times increasing as we reach over 6144 samples. There is also a moment where the MPPI-Generic library optimization time grows to be larger than the AutoRally implementation. That occurs when we switch from using the split kernels (Section III-C2) to the combined kernel (Section III-C3). The AutoRally implementation uses a combined kernel with fewer GPU synchronization points due to strictly requiring forward Euler integration for the dynamics. At the small hit to performance in the combined kernel, our library allows for many more features, such as multi-threaded cost functions, use of shared memory in the cost function, and implementation of more computationally-heavy integration methods such as Runge-Kutta or backward Euler integration. And while we see a hit to performance when using the combined kernel compared to AutoRally, we still see that the split kernel is faster for up to 2048 samples.
Figure 5: Optimization times for the TorchRL implementation across different CPUs and GPUs. TorchRL computation times are more dependent on the CPU as the RTX 3080 with an AMD 5600X ends up slower than a GTX 1050 Ti with an Intel 13900k.
The TorchRL implementation is notably performing quite poorly in Figs. 3 and 4 with runtimes being around $28$\mathrm{ms}$$ no matter the number of samples. Looking at TorchRL-specific results in Fig. 5, the TorchRL implementation seems to be heavily CPU-bound. A low-end GPU (1050 Ti) combined with a high-end CPU (Intel 13900K) can achieve better optimization times than a low-end CPU (AMD 5600X) combined with a high-end GPU (RTX 3080).
We also conducted tests on a Jetson Nano to show that even on relatively low-power and older systems, our library can still be used. As the latest version of CUDA supporting the Jetson Nano is 10.2 and the OS is Ubuntu 18.04, both the TorchRL and ROS2 MPPI implementations were not compatible. As such, we only have results comparing our MPPI-Generic implementation to the AutoRally implementation in Fig. 6. Here, the AutoRally implementation starts having faster compute times around 512 samples. Again, this is due to our library switching to the combined kernel which will be slower. However, our library on a Jetson Nano at 2048 samples has a roughly equivalent computation time to that of 2048 samples of the ROS2 implementation on the Intel 13900K process, showing that our GPU parallelization can allow for real-time optimization even on portable systems.
Figure 6: Optimization times for MPPI implementations on a Jetson Nano over a variety of number of samples. MPPI-Generic and AutoRally on this low-power hardware can still achieve sub-$10$\mathrm{ms}$$ optimization times for even 2048 samples. The AutoRally implementation quickly surpasses our implementation in optimization times.
In addition, we see the benefits of our library as we increase the computation time of the cost function. At this point, the TorchRL and ROS2 implementations have been shown to be slow in comparison to the other implementations and are thus dropped from this cost complexity comparison.
1float computeStateCost(…) {
2 float cost = 1.0;
3 for (int i = 0; i < NUM_COSINES; i++) {
4 cost = cos(cost);
5 }
6 cost *= 0.0;
7 // Continue to regular cost function
8
9}
List of Listings 14 Computation time inflation code added to the cost function. We add a configurable amount of calls to cos() as this is a computationally heavy function to run.
We artificially inflate the computation time of the cost function with Lst. 14 to judge how well the implementations scale to more complex cost functions. In Fig. 7, we see how increasing the computation time of the cost function scales for both implementations over the same hardware and for the same number of samples.
Figure 7: Optimization Times for MPPI-Generic and AutoRally implementations as the computation time of the cost function increases. Using an Intel 13900K, NVIDIA GTX 1050 Ti, and 8192 samples, our library implementation starts to outperform the AutoRally implementation when 20+ sequential cosine operations are added to the cost function.
V-B Comparisons to sampling-efficient algorithms
Figure 8: Average Accumulated Costs (left) and Optimization Times (right) with error bars signifying one standard deviation for a variety of step sizes and number of samples for DMD-MPC. $\times$ indicates the step size that achieves the lowest cost for a given number of samples. When using a low number of samples, a lower DMD-MPC step size provides the lowest average cost. However, as the number of samples increase, the best step size choice becomes $\gamma_{t}=1.0$ which is equivalent to the normal MPPI update law. With our library, increasing the number of samples to the point where the step size is no longer useful is still able to be run at over 800 Hz on the NVIDIA GTX 1050 Ti.
While we have shown that our implementation of MPPI can have faster computation times and a lot of flexibility in applications, there remains a question of how to balance between the number of samples and real-time performance. Our work decreases the computation time for sampling which in turn allows more samples in the same computation time, while other works have tried to reduce the amount of samples needed to evaluate the optimal control trajectory. Many authors [20, 21, 38, 42, 55, 56] have tried to do this by changing the sampling distribution; depending on the derivation of MPPI, different sampling distributions can be considered to have the same update rule. In [61], the authors introduce a new update rule through a generalization of MPC algorithms called Dynamic Mirror Descent Model Predictive Control (DMD-MPC) defined by the choice of shaping function, $S(\cdot)$, Sampling Distribution $\pi_{\theta}$, and Bregman Divergence $D_{\Psi}(\cdot,\cdot)$ which determines how close the new optimal control trajectory should remain to the previous. Using the exponential function, Gaussian sampling, and the KL Divergence, they derive a modification to the MPPI update law in Eq. 17 that introduces a step size parameter $\gamma_{t}\geq 0$:
${\bf u}_{t}^{k+1}=\left(1-\gamma_{t}\right){\bf u}_{t}^{k}+\gamma_{t}\frac{\mathbb{E}_{V\sim\pi_{\theta}}\left[\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right){\bf v}_{t}\right]}{\mathbb{E}_{V\sim\pi_{\theta}}\left[\exp\left(-\frac{1}{\lambda}\mathcal{J}\left(V\right)\right)\right]},$ (24)
where ${\bf v}_{t}$ is the control value at time $t$ from the sampled control sequence $V$, and $\mathcal{J}$ is as we defined in Eq. 3. In their results, the addition of a step size can improve performance when using a low number of samples. However, once the number of samples increases beyond a certain point, the optimal step size ends up being $1.0$ which is equivalent to the original MPPI update law. Having this option is useful in cases where you have a low computational budget. We show as long as you have a NVIDIA GPU from the last decade, you have enough computational budget to use more samples without needing to tune a step size.
Looking at Fig. 8, we ran a 2D double integrator system with state $[x,y,v_{x},v_{y}]$ and control $[a_{x},a_{y}]$ with the cost shown in Eq. 25:
$\begin{split}J&=1000\left(\text{1}_{\left\{\left(x^{2}+y^{2}\right)\leq 1.875^{2}\right\}}+\text{1}_{\left\{\left(x^{2}+y^{2}\right)\geq 2.125^{2}\right\}}\right)\\ &+2\left|2-\sqrt{v_{x}^{2}+v_{y}^{2}}\right|+2\left|4-\left(xv_{y}-yv_{x}\right)\right|.\end{split}$ (25)
[…]
VI Conclusion
In this paper, we introduce a new sampling-based optimization C++/CUDA library called MPPI-Generic. It contains implementations of MPPI, Tube-MPPI, and RMPPI controllers as well as an API that allows these controllers to be used with multiple dynamics and cost functions. We went through various ways that researchers could use this library, from using pre-defined dynamics and cost functions to implementing new sampling-based MPC controllers. We discussed the methods used to improve the computational performance and conducted performances comparisons against other widely-available implementations of MPPI over a variety of computer hardware to show the performance benefits our library can provide. Finally, we compared against a sample-efficient form of MPPI to show that with the speed improvements of our library, using more samples is a viable alternative with little hit to computation times. We plan to keep working on the library to add more capabilities and usage improvements such as a Python wrapper in the future.
TABLE III: MPPI Method Optimization Times at various number of samples on various hardware
CPU GPU Samples Method Avg. Time [$\mathrm{ms}$] CPU GPU Samples Method Avg. Time [$\mathrm{ms}$]
Intel 13900K 1060 6GB 128 MPPI-Generic 0.180 181±0.034 711 8uncertain0.1801810.03471180.180\,181$\pm$ 0.034\,711\,8start_ARG 0.180 181 end_ARG ± start_ARG 0.034 711 8 end_ARG Intel 13900K 1080 Ti 128 MPPI-Generic 0.179 371±0.028 955 6uncertain0.1793710.02895560.179\,371$\pm$ 0.028\,955\,6start_ARG 0.179 371 end_ARG ± start_ARG 0.028 955 6 end_ARG
Intel 13900K 1060 6GB 128 autorally 0.436 403±0.054 117 2uncertain0.4364030.05411720.436\,403$\pm$ 0.054\,117\,2start_ARG 0.436 403 end_ARG ± start_ARG 0.054 117 2 end_ARG Intel 13900K 1080 Ti 128 autorally 0.443 172±0.115 946uncertain0.4431720.1159460.443\,172$\pm$ 0.115\,946start_ARG 0.443 172 end_ARG ± start_ARG 0.115 946 end_ARG
Intel 13900K 1060 6GB 128 torchrl 24.556 195 97±1.644 165 462uncertain24.556195971.64416546224.556\,195\,97$\pm$ 1.644\,165\,462start_ARG 24.556 195 97 end_ARG ± start_ARG 1.644 165 462 end_ARG Intel 13900K 1080 Ti 128 torchrl 24.629 269 12±1.805 008 154uncertain24.629269121.80500815424.629\,269\,12$\pm$ 1.805\,008\,154start_ARG 24.629 269 12 end_ARG ± start_ARG 1.805 008 154 end_ARG
Intel 13900K 1060 6GB 256 MPPI-Generic 0.191 272±0.031 385uncertain0.1912720.0313850.191\,272$\pm$ 0.031\,385start_ARG 0.191 272 end_ARG ± start_ARG 0.031 385 end_ARG Intel 13900K 1080 Ti 256 MPPI-Generic 0.184 424±0.029 488 7uncertain0.1844240.02948870.184\,424$\pm$ 0.029\,488\,7start_ARG 0.184 424 end_ARG ± start_ARG 0.029 488 7 end_ARG
Intel 13900K 1060 6GB 256 autorally 0.431 935±0.037 263 4uncertain0.4319350.03726340.431\,935$\pm$ 0.037\,263\,4start_ARG 0.431 935 end_ARG ± start_ARG 0.037 263 4 end_ARG Intel 13900K 1080 Ti 256 autorally 0.438 665±0.035 979 7uncertain0.4386650.03597970.438\,665$\pm$ 0.035\,979\,7start_ARG 0.438 665 end_ARG ± start_ARG 0.035 979 7 end_ARG
Intel 13900K 1060 6GB 256 torchrl 28.415 566 68±1.401 850 333uncertain28.415566681.40185033328.415\,566\,68$\pm$ 1.401\,850\,333start_ARG 28.415 566 68 end_ARG ± start_ARG 1.401 850 333 end_ARG Intel 13900K 1080 Ti 256 torchrl 28.404 400 59±1.401 767 178uncertain28.404400591.40176717828.404\,400\,59$\pm$ 1.401\,767\,178start_ARG 28.404 400 59 end_ARG ± start_ARG 1.401 767 178 end_ARG
Intel 13900K 1060 6GB 512 MPPI-Generic 0.225 903±0.031 470 2uncertain0.2259030.03147020.225\,903$\pm$ 0.031\,470\,2start_ARG 0.225 903 end_ARG ± start_ARG 0.031 470 2 end_ARG Intel 13900K 1080 Ti 512 MPPI-Generic 0.200 678±0.030 628 6uncertain0.2006780.03062860.200\,678$\pm$ 0.030\,628\,6start_ARG 0.200 678 end_ARG ± start_ARG 0.030 628 6 end_ARG
Intel 13900K 1060 6GB 512 autorally 0.434 337±0.039 382 7uncertain0.4343370.03938270.434\,337$\pm$ 0.039\,382\,7start_ARG 0.434 337 end_ARG ± start_ARG 0.039 382 7 end_ARG Intel 13900K 1080 Ti 512 autorally 0.442 098±0.037 613 2uncertain0.4420980.03761320.442\,098$\pm$ 0.037\,613\,2start_ARG 0.442 098 end_ARG ± start_ARG 0.037 613 2 end_ARG
Intel 13900K 1060 6GB 512 torchrl 28.569 205 76±1.625 594 162uncertain28.569205761.62559416228.569\,205\,76$\pm$ 1.625\,594\,162start_ARG 28.569 205 76 end_ARG ± start_ARG 1.625 594 162 end_ARG Intel 13900K 1080 Ti 512 torchrl 28.445 260 76±1.600 661 798uncertain28.445260761.60066179828.445\,260\,76$\pm$ 1.600\,661\,798start_ARG 28.445 260 76 end_ARG ± start_ARG 1.600 661 798 end_ARG
Intel 13900K 1060 6GB 1024 MPPI-Generic 0.310 363±0.034 548 2uncertain0.3103630.03454820.310\,363$\pm$ 0.034\,548\,2start_ARG 0.310 363 end_ARG ± start_ARG 0.034 548 2 end_ARG Intel 13900K 1080 Ti 1024 MPPI-Generic 0.219 663±0.029 778 3uncertain0.2196630.02977830.219\,663$\pm$ 0.029\,778\,3start_ARG 0.219 663 end_ARG ± start_ARG 0.029 778 3 end_ARG
Intel 13900K 1060 6GB 1024 autorally 0.467 177±0.042 738 9uncertain0.4671770.04273890.467\,177$\pm$ 0.042\,738\,9start_ARG 0.467 177 end_ARG ± start_ARG 0.042 738 9 end_ARG Intel 13900K 1080 Ti 1024 autorally 0.447 832±0.037 942 3uncertain0.4478320.03794230.447\,832$\pm$ 0.037\,942\,3start_ARG 0.447 832 end_ARG ± start_ARG 0.037 942 3 end_ARG
Intel 13900K 1060 6GB 1024 torchrl 28.369 998 22±1.712 465 173uncertain28.369998221.71246517328.369\,998\,22$\pm$ 1.712\,465\,173start_ARG 28.369 998 22 end_ARG ± start_ARG 1.712 465 173 end_ARG Intel 13900K 1080 Ti 1024 torchrl 28.535 989 52±1.599 407 986uncertain28.535989521.59940798628.535\,989\,52$\pm$ 1.599\,407\,986start_ARG 28.535 989 52 end_ARG ± start_ARG 1.599 407 986 end_ARG
Intel 13900K 1060 6GB 2048 MPPI-Generic 0.479 197±0.039 354 7uncertain0.4791970.03935470.479\,197$\pm$ 0.039\,354\,7start_ARG 0.479 197 end_ARG ± start_ARG 0.039 354 7 end_ARG Intel 13900K 1080 Ti 2048 MPPI-Generic 0.297 774±0.033 314 9uncertain0.2977740.03331490.297\,774$\pm$ 0.033\,314\,9start_ARG 0.297 774 end_ARG ± start_ARG 0.033 314 9 end_ARG
Intel 13900K 1060 6GB 2048 autorally 0.544 643±0.067 192 6uncertain0.5446430.06719260.544\,643$\pm$ 0.067\,192\,6start_ARG 0.544 643 end_ARG ± start_ARG 0.067 192 6 end_ARG Intel 13900K 1080 Ti 2048 autorally 0.481 744±0.041 248 4uncertain0.4817440.04124840.481\,744$\pm$ 0.041\,248\,4start_ARG 0.481 744 end_ARG ± start_ARG 0.041 248 4 end_ARG
Intel 13900K 1060 6GB 2048 torchrl 28.592 346 43±1.469 444 317uncertain28.592346431.46944431728.592\,346\,43$\pm$ 1.469\,444\,317start_ARG 28.592 346 43 end_ARG ± start_ARG 1.469 444 317 end_ARG Intel 13900K 1080 Ti 2048 torchrl 28.657 792 09±1.390 036 982uncertain28.657792091.39003698228.657\,792\,09$\pm$ 1.390\,036\,982start_ARG 28.657 792 09 end_ARG ± start_ARG 1.390 036 982 end_ARG
Intel 13900K 1060 6GB 4096 MPPI-Generic 0.890 357±0.053 562 2uncertain0.8903570.05356220.890\,357$\pm$ 0.053\,562\,2start_ARG 0.890 357 end_ARG ± start_ARG 0.053 562 2 end_ARG Intel 13900K 1080 Ti 4096 MPPI-Generic 0.432 359±0.040 481 7uncertain0.4323590.04048170.432\,359$\pm$ 0.040\,481\,7start_ARG 0.432 359 end_ARG ± start_ARG 0.040 481 7 end_ARG
Intel 13900K 1060 6GB 4096 autorally 0.986 041±0.077 754 9uncertain0.9860410.07775490.986\,041$\pm$ 0.077\,754\,9start_ARG 0.986 041 end_ARG ± start_ARG 0.077 754 9 end_ARG Intel 13900K 1080 Ti 4096 autorally 0.548 638±0.042 029 9uncertain0.5486380.04202990.548\,638$\pm$ 0.042\,029\,9start_ARG 0.548 638 end_ARG ± start_ARG 0.042 029 9 end_ARG
Intel 13900K 1060 6GB 4096 torchrl 29.325 288 77±1.707 160 026uncertain29.325288771.70716002629.325\,288\,77$\pm$ 1.707\,160\,026start_ARG 29.325 288 77 end_ARG ± start_ARG 1.707 160 026 end_ARG Intel 13900K 1080 Ti 4096 torchrl 28.908 665 42±1.690 516 162uncertain28.908665421.69051616228.908\,665\,42$\pm$ 1.690\,516\,162start_ARG 28.908 665 42 end_ARG ± start_ARG 1.690 516 162 end_ARG
Intel 13900K 1060 6GB 6144 autorally 1.1427±0.077 404 7uncertain1.14270.07740471.1427$\pm$ 0.077\,404\,7start_ARG 1.1427 end_ARG ± start_ARG 0.077 404 7 end_ARG Intel 13900K 1080 Ti 6144 autorally 0.619 48±0.040 916 7uncertain0.619480.04091670.619\,48$\pm$ 0.040\,916\,7start_ARG 0.619 48 end_ARG ± start_ARG 0.040 916 7 end_ARG
Intel 13900K 1060 6GB 6144 MPPI-Generic 1.310 63±0.063 285 9uncertain1.310630.06328591.310\,63$\pm$ 0.063\,285\,9start_ARG 1.310 63 end_ARG ± start_ARG 0.063 285 9 end_ARG Intel 13900K 1080 Ti 6144 MPPI-Generic 0.701 594±0.045 963 1uncertain0.7015940.04596310.701\,594$\pm$ 0.045\,963\,1start_ARG 0.701 594 end_ARG ± start_ARG 0.045 963 1 end_ARG
Intel 13900K 1060 6GB 6144 torchrl 30.903 118 37±1.655 715 708uncertain30.903118371.65571570830.903\,118\,37$\pm$ 1.655\,715\,708start_ARG 30.903 118 37 end_ARG ± start_ARG 1.655 715 708 end_ARG Intel 13900K 1080 Ti 6144 torchrl 29.173 142 19±1.685 870 23uncertain29.173142191.6858702329.173\,142\,19$\pm$ 1.685\,870\,23start_ARG 29.173 142 19 end_ARG ± start_ARG 1.685 870 23 end_ARG
Intel 13900K 1060 6GB 8192 autorally 1.599 04±0.071 216 5uncertain1.599040.07121651.599\,04$\pm$ 0.071\,216\,5start_ARG 1.599 04 end_ARG ± start_ARG 0.071 216 5 end_ARG Intel 13900K 1080 Ti 8192 autorally 0.695 471±0.047 290 6uncertain0.6954710.04729060.695\,471$\pm$ 0.047\,290\,6start_ARG 0.695 471 end_ARG ± start_ARG 0.047 290 6 end_ARG
Intel 13900K 1060 6GB 8192 MPPI-Generic 1.755 41±0.070 203 4uncertain1.755410.07020341.755\,41$\pm$ 0.070\,203\,4start_ARG 1.755 41 end_ARG ± start_ARG 0.070 203 4 end_ARG Intel 13900K 1080 Ti 8192 MPPI-Generic 1.109 35±0.057 626 7uncertain1.109350.05762671.109\,35$\pm$ 0.057\,626\,7start_ARG 1.109 35 end_ARG ± start_ARG 0.057 626 7 end_ARG
Intel 13900K 1060 6GB 8192 torchrl 31.716 877 22±1.393 955 124uncertain31.716877221.39395512431.716\,877\,22$\pm$ 1.393\,955\,124start_ARG 31.716 877 22 end_ARG ± start_ARG 1.393 955 124 end_ARG Intel 13900K 1080 Ti 8192 torchrl 29.743 070 36±1.505 358 741uncertain29.743070361.50535874129.743\,070\,36$\pm$ 1.505\,358\,741start_ARG 29.743 070 36 end_ARG ± start_ARG 1.505 358 741 end_ARG
Intel 13900K 1060 6GB 16384 autorally 3.045 68±0.635 748uncertain3.045680.6357483.045\,68$\pm$ 0.635\,748start_ARG 3.045 68 end_ARG ± start_ARG 0.635 748 end_ARG Intel 13900K 1080 Ti 16384 autorally 1.334 03±0.060 558 9uncertain1.334030.06055891.334\,03$\pm$ 0.060\,558\,9start_ARG 1.334 03 end_ARG ± start_ARG 0.060 558 9 end_ARG
Intel 13900K 1060 6GB 16384 MPPI-Generic 3.337 23±0.098 572uncertain3.337230.0985723.337\,23$\pm$ 0.098\,572start_ARG 3.337 23 end_ARG ± start_ARG 0.098 572 end_ARG Intel 13900K 1080 Ti 16384 MPPI-Generic 2.187 85±0.078 458 1uncertain2.187850.07845812.187\,85$\pm$ 0.078\,458\,1start_ARG 2.187 85 end_ARG ± start_ARG 0.078 458 1 end_ARG
Intel 13900K 1060 6GB 16384 torchrl 41.260 504 01±1.609 524 793uncertain41.260504011.60952479341.260\,504\,01$\pm$ 1.609\,524\,793start_ARG 41.260 504 01 end_ARG ± start_ARG 1.609 524 793 end_ARG Intel 13900K 1080 Ti 16384 torchrl 33.756 266 12±1.568 212 997uncertain33.756266121.56821299733.756\,266\,12$\pm$ 1.568\,212\,997start_ARG 33.756 266 12 end_ARG ± start_ARG 1.568 212 997 end_ARG
Intel 13900K 2080 128 MPPI-Generic 0.165 901±0.023 169 1uncertain0.1659010.02316910.165\,901$\pm$ 0.023\,169\,1start_ARG 0.165 901 end_ARG ± start_ARG 0.023 169 1 end_ARG Intel 13900K 4090 128 MPPI-Generic 0.122 728±0.031 763 5uncertain0.1227280.03176350.122\,728$\pm$ 0.031\,763\,5start_ARG 0.122 728 end_ARG ± start_ARG 0.031 763 5 end_ARG
Intel 13900K 2080 128 autorally 0.293 212±0.137 769uncertain0.2932120.1377690.293\,212$\pm$ 0.137\,769start_ARG 0.293 212 end_ARG ± start_ARG 0.137 769 end_ARG Intel 13900K 4090 128 autorally 0.197 862±0.038 947 1uncertain0.1978620.03894710.197\,862$\pm$ 0.038\,947\,1start_ARG 0.197 862 end_ARG ± start_ARG 0.038 947 1 end_ARG
Intel 13900K 2080 128 torchrl 24.817 154 65±1.604 587 402uncertain24.817154651.60458740224.817\,154\,65$\pm$ 1.604\,587\,402start_ARG 24.817 154 65 end_ARG ± start_ARG 1.604 587 402 end_ARG Intel 13900K 4090 128 torchrl 24.471 596 96±1.629 851 147uncertain24.471596961.62985114724.471\,596\,96$\pm$ 1.629\,851\,147start_ARG 24.471 596 96 end_ARG ± start_ARG 1.629 851 147 end_ARG
Intel 13900K 2080 256 MPPI-Generic 0.171 244±0.024 012 6uncertain0.1712440.02401260.171\,244$\pm$ 0.024\,012\,6start_ARG 0.171 244 end_ARG ± start_ARG 0.024 012 6 end_ARG Intel 13900K 4090 256 MPPI-Generic 0.121 906±0.028 788 8uncertain0.1219060.02878880.121\,906$\pm$ 0.028\,788\,8start_ARG 0.121 906 end_ARG ± start_ARG 0.028 788 8 end_ARG
Intel 13900K 2080 256 autorally 0.291 783±0.028 158 7uncertain0.2917830.02815870.291\,783$\pm$ 0.028\,158\,7start_ARG 0.291 783 end_ARG ± start_ARG 0.028 158 7 end_ARG Intel 13900K 4090 256 autorally 0.196 282±0.029 150 2uncertain0.1962820.02915020.196\,282$\pm$ 0.029\,150\,2start_ARG 0.196 282 end_ARG ± start_ARG 0.029 150 2 end_ARG
Intel 13900K 2080 256 torchrl 28.632 213 35±1.294 271 899uncertain28.632213351.29427189928.632\,213\,35$\pm$ 1.294\,271\,899start_ARG 28.632 213 35 end_ARG ± start_ARG 1.294 271 899 end_ARG Intel 13900K 4090 256 torchrl 28.166 282 42±1.328 607 151uncertain28.166282421.32860715128.166\,282\,42$\pm$ 1.328\,607\,151start_ARG 28.166 282 42 end_ARG ± start_ARG 1.328 607 151 end_ARG
Intel 13900K 2080 512 MPPI-Generic 0.180 227±0.028 030 5uncertain0.1802270.02803050.180\,227$\pm$ 0.028\,030\,5start_ARG 0.180 227 end_ARG ± start_ARG 0.028 030 5 end_ARG Intel 13900K 4090 512 MPPI-Generic 0.123 621±0.026 492 9uncertain0.1236210.02649290.123\,621$\pm$ 0.026\,492\,9start_ARG 0.123 621 end_ARG ± start_ARG 0.026 492 9 end_ARG
Intel 13900K 2080 512 autorally 0.299 438±0.031 532 8uncertain0.2994380.03153280.299\,438$\pm$ 0.031\,532\,8start_ARG 0.299 438 end_ARG ± start_ARG 0.031 532 8 end_ARG Intel 13900K 4090 512 autorally 0.197 492±0.028 059 2uncertain0.1974920.02805920.197\,492$\pm$ 0.028\,059\,2start_ARG 0.197 492 end_ARG ± start_ARG 0.028 059 2 end_ARG
Intel 13900K 2080 512 torchrl 28.703 887 94±1.545 536 58uncertain28.703887941.5455365828.703\,887\,94$\pm$ 1.545\,536\,58start_ARG 28.703 887 94 end_ARG ± start_ARG 1.545 536 58 end_ARG Intel 13900K 4090 512 torchrl 28.220 973 49±1.604 033 825uncertain28.220973491.60403382528.220\,973\,49$\pm$ 1.604\,033\,825start_ARG 28.220 973 49 end_ARG ± start_ARG 1.604 033 825 end_ARG
Intel 13900K 2080 1024 MPPI-Generic 0.203 835±0.027 832 3uncertain0.2038350.02783230.203\,835$\pm$ 0.027\,832\,3start_ARG 0.203 835 end_ARG ± start_ARG 0.027 832 3 end_ARG Intel 13900K 4090 1024 MPPI-Generic 0.130 65±0.025 213 3uncertain0.130650.02521330.130\,65$\pm$ 0.025\,213\,3start_ARG 0.130 65 end_ARG ± start_ARG 0.025 213 3 end_ARG
Intel 13900K 2080 1024 autorally 0.302 238±0.028 387uncertain0.3022380.0283870.302\,238$\pm$ 0.028\,387start_ARG 0.302 238 end_ARG ± start_ARG 0.028 387 end_ARG Intel 13900K 4090 1024 autorally 0.202 446±0.024 600 8uncertain0.2024460.02460080.202\,446$\pm$ 0.024\,600\,8start_ARG 0.202 446 end_ARG ± start_ARG 0.024 600 8 end_ARG
Intel 13900K 2080 1024 torchrl 28.763 414 62±1.535 270 445uncertain28.763414621.53527044528.763\,414\,62$\pm$ 1.535\,270\,445start_ARG 28.763 414 62 end_ARG ± start_ARG 1.535 270 445 end_ARG Intel 13900K 4090 1024 torchrl 28.136 604 31±1.556 810 549uncertain28.136604311.55681054928.136\,604\,31$\pm$ 1.556\,810\,549start_ARG 28.136 604 31 end_ARG ± start_ARG 1.556 810 549 end_ARG
Intel 13900K 2080 2048 MPPI-Generic 0.260 856±0.026 15uncertain0.2608560.026150.260\,856$\pm$ 0.026\,15start_ARG 0.260 856 end_ARG ± start_ARG 0.026 15 end_ARG Intel 13900K 4090 2048 MPPI-Generic 0.146 179±0.036 207 3uncertain0.1461790.03620730.146\,179$\pm$ 0.036\,207\,3start_ARG 0.146 179 end_ARG ± start_ARG 0.036 207 3 end_ARG
Intel 13900K 2080 2048 autorally 0.322 63±0.033 309 4uncertain0.322630.03330940.322\,63$\pm$ 0.033\,309\,4start_ARG 0.322 63 end_ARG ± start_ARG 0.033 309 4 end_ARG Intel 13900K 4090 2048 autorally 0.213 538±0.034 578 4uncertain0.2135380.03457840.213\,538$\pm$ 0.034\,578\,4start_ARG 0.213 538 end_ARG ± start_ARG 0.034 578 4 end_ARG
Intel 13900K 2080 2048 torchrl 28.826 831 58±1.342 341 806uncertain28.826831581.34234180628.826\,831\,58$\pm$ 1.342\,341\,806start_ARG 28.826 831 58 end_ARG ± start_ARG 1.342 341 806 end_ARG Intel 13900K 4090 2048 torchrl 28.205 497 26±1.322 743 723uncertain28.205497261.32274372328.205\,497\,26$\pm$ 1.322\,743\,723start_ARG 28.205 497 26 end_ARG ± start_ARG 1.322 743 723 end_ARG
Intel 13900K 2080 4096 MPPI-Generic 0.397 538±0.034 400 2uncertain0.3975380.03440020.397\,538$\pm$ 0.034\,400\,2start_ARG 0.397 538 end_ARG ± start_ARG 0.034 400 2 end_ARG Intel 13900K 4090 4096 MPPI-Generic 0.169 998±0.029 523 2uncertain0.1699980.02952320.169\,998$\pm$ 0.029\,523\,2start_ARG 0.169 998 end_ARG ± start_ARG 0.029 523 2 end_ARG
Intel 13900K 2080 4096 autorally 0.403 02±0.034 771 3uncertain0.403020.03477130.403\,02$\pm$ 0.034\,771\,3start_ARG 0.403 02 end_ARG ± start_ARG 0.034 771 3 end_ARG Intel 13900K 4090 4096 autorally 0.224 06±0.028 372 3uncertain0.224060.02837230.224\,06$\pm$ 0.028\,372\,3start_ARG 0.224 06 end_ARG ± start_ARG 0.028 372 3 end_ARG
Intel 13900K 2080 4096 torchrl 29.099 217 18±1.590 309 02uncertain29.099217181.5903090229.099\,217\,18$\pm$ 1.590\,309\,02start_ARG 29.099 217 18 end_ARG ± start_ARG 1.590 309 02 end_ARG Intel 13900K 4090 4096 torchrl 28.245 249 51±1.561 157 368uncertain28.245249511.56115736828.245\,249\,51$\pm$ 1.561\,157\,368start_ARG 28.245 249 51 end_ARG ± start_ARG 1.561 157 368 end_ARG
Intel 13900K 2080 6144 autorally 0.518 005±0.039 834 1uncertain0.5180050.03983410.518\,005$\pm$ 0.039\,834\,1start_ARG 0.518 005 end_ARG ± start_ARG 0.039 834 1 end_ARG Intel 13900K 4090 6144 MPPI-Generic 0.200 644±0.027 676 1uncertain0.2006440.02767610.200\,644$\pm$ 0.027\,676\,1start_ARG 0.200 644 end_ARG ± start_ARG 0.027 676 1 end_ARG
Intel 13900K 2080 6144 MPPI-Generic 0.576 165±0.042 399 6uncertain0.5761650.04239960.576\,165$\pm$ 0.042\,399\,6start_ARG 0.576 165 end_ARG ± start_ARG 0.042 399 6 end_ARG Intel 13900K 4090 6144 autorally 0.232 756±0.032 101 9uncertain0.2327560.03210190.232\,756$\pm$ 0.032\,101\,9start_ARG 0.232 756 end_ARG ± start_ARG 0.032 101 9 end_ARG
Intel 13900K 2080 6144 torchrl 29.138 362 65±1.585 337 189uncertain29.138362651.58533718929.138\,362\,65$\pm$ 1.585\,337\,189start_ARG 29.138 362 65 end_ARG ± start_ARG 1.585 337 189 end_ARG Intel 13900K 4090 6144 torchrl 28.357 528 92±1.587 800 846uncertain28.357528921.58780084628.357\,528\,92$\pm$ 1.587\,800\,846start_ARG 28.357 528 92 end_ARG ± start_ARG 1.587 800 846 end_ARG
Intel 13900K 2080 8192 autorally 0.580 367±0.043 674 8uncertain0.5803670.04367480.580\,367$\pm$ 0.043\,674\,8start_ARG 0.580 367 end_ARG ± start_ARG 0.043 674 8 end_ARG Intel 13900K 4090 8192 MPPI-Generic 0.231 558±0.029 574 6uncertain0.2315580.02957460.231\,558$\pm$ 0.029\,574\,6start_ARG 0.231 558 end_ARG ± start_ARG 0.029 574 6 end_ARG
Intel 13900K 2080 8192 MPPI-Generic 0.704 521±0.042 760 7uncertain0.7045210.04276070.704\,521$\pm$ 0.042\,760\,7start_ARG 0.704 521 end_ARG ± start_ARG 0.042 760 7 end_ARG Intel 13900K 4090 8192 autorally 0.245 015±0.026 584 6uncertain0.2450150.02658460.245\,015$\pm$ 0.026\,584\,6start_ARG 0.245 015 end_ARG ± start_ARG 0.026 584 6 end_ARG
Intel 13900K 2080 8192 torchrl 29.417 226 79±1.276 422 812uncertain29.417226791.27642281229.417\,226\,79$\pm$ 1.276\,422\,812start_ARG 29.417 226 79 end_ARG ± start_ARG 1.276 422 812 end_ARG Intel 13900K 4090 8192 torchrl 28.590 681 31±1.507 338 15uncertain28.590681311.5073381528.590\,681\,31$\pm$ 1.507\,338\,15start_ARG 28.590 681 31 end_ARG ± start_ARG 1.507 338 15 end_ARG
Intel 13900K 2080 16384 autorally 1.192 45±0.059 349 6uncertain1.192450.05934961.192\,45$\pm$ 0.059\,349\,6start_ARG 1.192 45 end_ARG ± start_ARG 0.059 349 6 end_ARG Intel 13900K 4090 16384 autorally 0.338 403±0.032 683 6uncertain0.3384030.03268360.338\,403$\pm$ 0.032\,683\,6start_ARG 0.338 403 end_ARG ± start_ARG 0.032 683 6 end_ARG
Intel 13900K 2080 16384 MPPI-Generic 1.418 27±0.061 324 2uncertain1.418270.06132421.418\,27$\pm$ 0.061\,324\,2start_ARG 1.418 27 end_ARG ± start_ARG 0.061 324 2 end_ARG Intel 13900K 4090 16384 MPPI-Generic 0.388 618±0.038 214uncertain0.3886180.0382140.388\,618$\pm$ 0.038\,214start_ARG 0.388 618 end_ARG ± start_ARG 0.038 214 end_ARG
Intel 13900K 2080 16384 torchrl 32.369 725 47±1.529 793 824uncertain32.369725471.52979382432.369\,725\,47$\pm$ 1.529\,793\,824start_ARG 32.369 725 47 end_ARG ± start_ARG 1.529 793 824 end_ARG Intel 13900K 4090 16384 torchrl 28.604 603 29±1.627 474 615uncertain28.604603291.62747461528.604\,603\,29$\pm$ 1.627\,474\,615start_ARG 28.604 603 29 end_ARG ± start_ARG 1.627 474 615 end_ARG
CPU GPU Samples Method Avg. Time [$\mathrm{ms}$] CPU GPU Samples Method Avg. Time [$\mathrm{ms}$]
AMD 5600X 3080 128 MPPI-Generic 0.155 392±0.015 599 4uncertain0.1553920.01559940.155\,392$\pm$ 0.015\,599\,4start_ARG 0.155 392 end_ARG ± start_ARG 0.015 599 4 end_ARG Intel 13900K 3080 128 MPPI-Generic 0.147 546±0.031 614 2uncertain0.1475460.03161420.147\,546$\pm$ 0.031\,614\,2start_ARG 0.147 546 end_ARG ± start_ARG 0.031 614 2 end_ARG
AMD 5600X 3080 128 autorally 0.270 118±0.016 916 2uncertain0.2701180.01691620.270\,118$\pm$ 0.016\,916\,2start_ARG 0.270 118 end_ARG ± start_ARG 0.016 916 2 end_ARG Intel 13900K 3080 128 autorally 0.245 459±0.118 456uncertain0.2454590.1184560.245\,459$\pm$ 0.118\,456start_ARG 0.245 459 end_ARG ± start_ARG 0.118 456 end_ARG
AMD 5600X 128 ros2 0.615 263±0.031 906 2uncertain0.6152630.03190620.615\,263$\pm$ 0.031\,906\,2start_ARG 0.615 263 end_ARG ± start_ARG 0.031 906 2 end_ARG Intel 13900K 128 ros2 0.389 923±0.016 247 6uncertain0.3899230.01624760.389\,923$\pm$ 0.016\,247\,6start_ARG 0.389 923 end_ARG ± start_ARG 0.016 247 6 end_ARG
AMD 5600X 3080 128 torchrl 47.610 795 02±1.583 180 632uncertain47.610795021.58318063247.610\,795\,02$\pm$ 1.583\,180\,632start_ARG 47.610 795 02 end_ARG ± start_ARG 1.583 180 632 end_ARG Intel 13900K 3080 128 torchrl 24.774 477 72±1.728 524 408uncertain24.774477721.72852440824.774\,477\,72$\pm$ 1.728\,524\,408start_ARG 24.774 477 72 end_ARG ± start_ARG 1.728 524 408 end_ARG
AMD 5600X 3080 256 MPPI-Generic 0.153 569±0.033 578 4uncertain0.1535690.03357840.153\,569$\pm$ 0.033\,578\,4start_ARG 0.153 569 end_ARG ± start_ARG 0.033 578 4 end_ARG Intel 13900K 3080 256 MPPI-Generic 0.152 581±0.027 063uncertain0.1525810.0270630.152\,581$\pm$ 0.027\,063start_ARG 0.152 581 end_ARG ± start_ARG 0.027 063 end_ARG
AMD 5600X 3080 256 autorally 0.274 294±0.015 156 5uncertain0.2742940.01515650.274\,294$\pm$ 0.015\,156\,5start_ARG 0.274 294 end_ARG ± start_ARG 0.015 156 5 end_ARG Intel 13900K 3080 256 autorally 0.242 785±0.029 394uncertain0.2427850.0293940.242\,785$\pm$ 0.029\,394start_ARG 0.242 785 end_ARG ± start_ARG 0.029 394 end_ARG
AMD 5600X 256 ros2 1.080 37±0.042 655 8uncertain1.080370.04265581.080\,37$\pm$ 0.042\,655\,8start_ARG 1.080 37 end_ARG ± start_ARG 0.042 655 8 end_ARG Intel 13900K 256 ros2 0.728 985±0.010 112 1uncertain0.7289850.01011210.728\,985$\pm$ 0.010\,112\,1start_ARG 0.728 985 end_ARG ± start_ARG 0.010 112 1 end_ARG
AMD 5600X 3080 256 torchrl 53.877 165 32±2.024 329 58uncertain53.877165322.0243295853.877\,165\,32$\pm$ 2.024\,329\,58start_ARG 53.877 165 32 end_ARG ± start_ARG 2.024 329 58 end_ARG Intel 13900K 3080 256 torchrl 28.486 258 75±1.450 323 227uncertain28.486258751.45032322728.486\,258\,75$\pm$ 1.450\,323\,227start_ARG 28.486 258 75 end_ARG ± start_ARG 1.450 323 227 end_ARG
AMD 5600X 3080 512 MPPI-Generic 0.152 786±0.011 513 3uncertain0.1527860.01151330.152\,786$\pm$ 0.011\,513\,3start_ARG 0.152 786 end_ARG ± start_ARG 0.011 513 3 end_ARG Intel 13900K 3080 512 MPPI-Generic 0.153 557±0.034 477 8uncertain0.1535570.03447780.153\,557$\pm$ 0.034\,477\,8start_ARG 0.153 557 end_ARG ± start_ARG 0.034 477 8 end_ARG
AMD 5600X 3080 512 autorally 0.269 58±0.012 344 5uncertain0.269580.01234450.269\,58$\pm$ 0.012\,344\,5start_ARG 0.269 58 end_ARG ± start_ARG 0.012 344 5 end_ARG Intel 13900K 3080 512 autorally 0.246 625±0.036 877 7uncertain0.2466250.03687770.246\,625$\pm$ 0.036\,877\,7start_ARG 0.246 625 end_ARG ± start_ARG 0.036 877 7 end_ARG
AMD 5600X 512 ros2 2.101 14±0.096 183 5uncertain2.101140.09618352.101\,14$\pm$ 0.096\,183\,5start_ARG 2.101 14 end_ARG ± start_ARG 0.096 183 5 end_ARG Intel 13900K 512 ros2 1.418 62±0.044 525 8uncertain1.418620.04452581.418\,62$\pm$ 0.044\,525\,8start_ARG 1.418 62 end_ARG ± start_ARG 0.044 525 8 end_ARG
AMD 5600X 3080 512 torchrl 53.889 682 77±2.021 575 269uncertain53.889682772.02157526953.889\,682\,77$\pm$ 2.021\,575\,269start_ARG 53.889 682 77 end_ARG ± start_ARG 2.021 575 269 end_ARG Intel 13900K 3080 512 torchrl 28.583 495 38±1.653 015 595uncertain28.583495381.65301559528.583\,495\,38$\pm$ 1.653\,015\,595start_ARG 28.583 495 38 end_ARG ± start_ARG 1.653 015 595 end_ARG
AMD 5600X 3080 1024 MPPI-Generic 0.170 408±0.015 452 6uncertain0.1704080.01545260.170\,408$\pm$ 0.015\,452\,6start_ARG 0.170 408 end_ARG ± start_ARG 0.015 452 6 end_ARG Intel 13900K 3080 1024 MPPI-Generic 0.163 43±0.027 135 2uncertain0.163430.02713520.163\,43$\pm$ 0.027\,135\,2start_ARG 0.163 43 end_ARG ± start_ARG 0.027 135 2 end_ARG
AMD 5600X 3080 1024 autorally 0.275 659±0.014 764 3uncertain0.2756590.01476430.275\,659$\pm$ 0.014\,764\,3start_ARG 0.275 659 end_ARG ± start_ARG 0.014 764 3 end_ARG Intel 13900K 3080 1024 autorally 0.259 898±0.035 735 1uncertain0.2598980.03573510.259\,898$\pm$ 0.035\,735\,1start_ARG 0.259 898 end_ARG ± start_ARG 0.035 735 1 end_ARG
AMD 5600X 1024 ros2 4.751 78±0.240 294uncertain4.751780.2402944.751\,78$\pm$ 0.240\,294start_ARG 4.751 78 end_ARG ± start_ARG 0.240 294 end_ARG Intel 13900K 1024 ros2 3.355 86±0.035 017 1uncertain3.355860.03501713.355\,86$\pm$ 0.035\,017\,1start_ARG 3.355 86 end_ARG ± start_ARG 0.035 017 1 end_ARG
AMD 5600X 3080 1024 torchrl 53.965 529 68±1.802 219 12uncertain53.965529681.8022191253.965\,529\,68$\pm$ 1.802\,219\,12start_ARG 53.965 529 68 end_ARG ± start_ARG 1.802 219 12 end_ARG Intel 13900K 3080 1024 torchrl 28.677 164 79±1.746 510 696uncertain28.677164791.74651069628.677\,164\,79$\pm$ 1.746\,510\,696start_ARG 28.677 164 79 end_ARG ± start_ARG 1.746 510 696 end_ARG
AMD 5600X 3080 2048 MPPI-Generic 0.197 253±0.013 828uncertain0.1972530.0138280.197\,253$\pm$ 0.013\,828start_ARG 0.197 253 end_ARG ± start_ARG 0.013 828 end_ARG Intel 13900K 3080 2048 MPPI-Generic 0.191 848±0.033 05uncertain0.1918480.033050.191\,848$\pm$ 0.033\,05start_ARG 0.191 848 end_ARG ± start_ARG 0.033 05 end_ARG
AMD 5600X 3080 2048 autorally 0.285 042±0.014 258 8uncertain0.2850420.01425880.285\,042$\pm$ 0.014\,258\,8start_ARG 0.285 042 end_ARG ± start_ARG 0.014 258 8 end_ARG Intel 13900K 3080 2048 autorally 0.262 051±0.038 966 9uncertain0.2620510.03896690.262\,051$\pm$ 0.038\,966\,9start_ARG 0.262 051 end_ARG ± start_ARG 0.038 966 9 end_ARG
AMD 5600X 2048 ros2 9.499 57±0.379 606uncertain9.499570.3796069.499\,57$\pm$ 0.379\,606start_ARG 9.499 57 end_ARG ± start_ARG 0.379 606 end_ARG Intel 13900K 2048 ros2 6.721 17±0.065 063 8uncertain6.721170.06506386.721\,17$\pm$ 0.065\,063\,8start_ARG 6.721 17 end_ARG ± start_ARG 0.065 063 8 end_ARG
AMD 5600X 3080 2048 torchrl 54.020 420 31±2.001 666 66uncertain54.020420312.0016666654.020\,420\,31$\pm$ 2.001\,666\,66start_ARG 54.020 420 31 end_ARG ± start_ARG 2.001 666 66 end_ARG Intel 13900K 3080 2048 torchrl 28.675 570 49±1.463 939 992uncertain28.675570491.46393999228.675\,570\,49$\pm$ 1.463\,939\,992start_ARG 28.675 570 49 end_ARG ± start_ARG 1.463 939 992 end_ARG
AMD 5600X 3080 4096 MPPI-Generic 0.256 218±0.037 690 8uncertain0.2562180.03769080.256\,218$\pm$ 0.037\,690\,8start_ARG 0.256 218 end_ARG ± start_ARG 0.037 690 8 end_ARG Intel 13900K 3080 4096 MPPI-Generic 0.248 872±0.038 845 7uncertain0.2488720.03884570.248\,872$\pm$ 0.038\,845\,7start_ARG 0.248 872 end_ARG ± start_ARG 0.038 845 7 end_ARG
AMD 5600X 3080 4096 autorally 0.289 591±0.014 270 1uncertain0.2895910.01427010.289\,591$\pm$ 0.014\,270\,1start_ARG 0.289 591 end_ARG ± start_ARG 0.014 270 1 end_ARG Intel 13900K 3080 4096 autorally 0.275 967±0.042 559 5uncertain0.2759670.04255950.275\,967$\pm$ 0.042\,559\,5start_ARG 0.275 967 end_ARG ± start_ARG 0.042 559 5 end_ARG
AMD 5600X 4096 ros2 19.5829±0.788 82uncertain19.58290.7888219.5829$\pm$ 0.788\,82start_ARG 19.5829 end_ARG ± start_ARG 0.788 82 end_ARG Intel 13900K 4096 ros2 13.9169±0.151 997uncertain13.91690.15199713.9169$\pm$ 0.151\,997start_ARG 13.9169 end_ARG ± start_ARG 0.151 997 end_ARG
AMD 5600X 3080 4096 torchrl 54.148 410 08±2.084 040 739uncertain54.148410082.08404073954.148\,410\,08$\pm$ 2.084\,040\,739start_ARG 54.148 410 08 end_ARG ± start_ARG 2.084 040 739 end_ARG Intel 13900K 3080 4096 torchrl 28.756 500 24±1.631 511 033uncertain28.756500241.63151103328.756\,500\,24$\pm$ 1.631\,511\,033start_ARG 28.756 500 24 end_ARG ± start_ARG 1.631 511 033 end_ARG
AMD 5600X 3080 6144 MPPI-Generic 0.314 502±0.019 407 6uncertain0.3145020.01940760.314\,502$\pm$ 0.019\,407\,6start_ARG 0.314 502 end_ARG ± start_ARG 0.019 407 6 end_ARG Intel 13900K 3080 6144 MPPI-Generic 0.306 363±0.039 101 6uncertain0.3063630.03910160.306\,363$\pm$ 0.039\,101\,6start_ARG 0.306 363 end_ARG ± start_ARG 0.039 101 6 end_ARG
AMD 5600X 3080 6144 autorally 0.336 179±0.014 264 1uncertain0.3361790.01426410.336\,179$\pm$ 0.014\,264\,1start_ARG 0.336 179 end_ARG ± start_ARG 0.014 264 1 end_ARG Intel 13900K 3080 6144 autorally 0.324 317±0.043 043 3uncertain0.3243170.04304330.324\,317$\pm$ 0.043\,043\,3start_ARG 0.324 317 end_ARG ± start_ARG 0.043 043 3 end_ARG
AMD 5600X 6144 ros2 29.8719±0.864 029uncertain29.87190.86402929.8719$\pm$ 0.864\,029start_ARG 29.8719 end_ARG ± start_ARG 0.864 029 end_ARG Intel 13900K 6144 ros2 21.7392±0.188 897uncertain21.73920.18889721.7392$\pm$ 0.188\,897start_ARG 21.7392 end_ARG ± start_ARG 0.188 897 end_ARG
AMD 5600X 3080 6144 torchrl 54.183 118 58±1.787 616 743uncertain54.183118581.78761674354.183\,118\,58$\pm$ 1.787\,616\,743start_ARG 54.183 118 58 end_ARG ± start_ARG 1.787 616 743 end_ARG Intel 13900K 3080 6144 torchrl 28.830 320 84±1.647 170 024uncertain28.830320841.64717002428.830\,320\,84$\pm$ 1.647\,170\,024start_ARG 28.830 320 84 end_ARG ± start_ARG 1.647 170 024 end_ARG
AMD 5600X 3080 8192 autorally 0.366 365±0.016 713 6uncertain0.3663650.01671360.366\,365$\pm$ 0.016\,713\,6start_ARG 0.366 365 end_ARG ± start_ARG 0.016 713 6 end_ARG Intel 13900K 3080 8192 MPPI-Generic 0.358 96±0.037 131 9uncertain0.358960.03713190.358\,96$\pm$ 0.037\,131\,9start_ARG 0.358 96 end_ARG ± start_ARG 0.037 131 9 end_ARG
AMD 5600X 3080 8192 MPPI-Generic 0.369 283±0.021 046 5uncertain0.3692830.02104650.369\,283$\pm$ 0.021\,046\,5start_ARG 0.369 283 end_ARG ± start_ARG 0.021 046 5 end_ARG Intel 13900K 3080 8192 autorally 0.360 461±0.044 851 2uncertain0.3604610.04485120.360\,461$\pm$ 0.044\,851\,2start_ARG 0.360 461 end_ARG ± start_ARG 0.044 851 2 end_ARG
AMD 5600X 8192 ros2 40.6365±1.059 23uncertain40.63651.0592340.6365$\pm$ 1.059\,23start_ARG 40.6365 end_ARG ± start_ARG 1.059 23 end_ARG Intel 13900K 3080 8192 torchrl 28.978 863±1.443 538 958uncertain28.9788631.44353895828.978\,863$\pm$ 1.443\,538\,958start_ARG 28.978 863 end_ARG ± start_ARG 1.443 538 958 end_ARG
AMD 5600X 3080 8192 torchrl 54.198 904 75±2.093 610 932uncertain54.198904752.09361093254.198\,904\,75$\pm$ 2.093\,610\,932start_ARG 54.198 904 75 end_ARG ± start_ARG 2.093 610 932 end_ARG Intel 13900K 8192 ros2 29.9169±0.239 302uncertain29.91690.23930229.9169$\pm$ 0.239\,302start_ARG 29.9169 end_ARG ± start_ARG 0.239 302 end_ARG
AMD 5600X 3080 16384 autorally 0.512 088±0.018 260 1uncertain0.5120880.01826010.512\,088$\pm$ 0.018\,260\,1start_ARG 0.512 088 end_ARG ± start_ARG 0.018 260 1 end_ARG Intel 13900K 3080 16384 autorally 0.499 134±0.049 703 4uncertain0.4991340.04970340.499\,134$\pm$ 0.049\,703\,4start_ARG 0.499 134 end_ARG ± start_ARG 0.049 703 4 end_ARG
AMD 5600X 3080 16384 MPPI-Generic 0.578 981±0.024 072 7uncertain0.5789810.02407270.578\,981$\pm$ 0.024\,072\,7start_ARG 0.578 981 end_ARG ± start_ARG 0.024 072 7 end_ARG Intel 13900K 3080 16384 MPPI-Generic 0.627 083±0.047 484 6uncertain0.6270830.04748460.627\,083$\pm$ 0.047\,484\,6start_ARG 0.627 083 end_ARG ± start_ARG 0.047 484 6 end_ARG
AMD 5600X 3080 16384 torchrl 54.679 183 01±1.695 322 846uncertain54.679183011.69532284654.679\,183\,01$\pm$ 1.695\,322\,846start_ARG 54.679 183 01 end_ARG ± start_ARG 1.695 322 846 end_ARG Intel 13900K 3080 16384 torchrl 29.977 583 17±1.674 777 268uncertain29.977583171.67477726829.977\,583\,17$\pm$ 1.674\,777\,268start_ARG 29.977 583 17 end_ARG ± start_ARG 1.674 777 268 end_ARG
AMD 5600X 16384 ros2 84.2601±1.646 13uncertain84.26011.6461384.2601$\pm$ 1.646\,13start_ARG 84.2601 end_ARG ± start_ARG 1.646 13 end_ARG Intel 13900K 16384 ros2 62.2642±0.990 145uncertain62.26420.99014562.2642$\pm$ 0.990\,145start_ARG 62.2642 end_ARG ± start_ARG 0.990 145 end_ARG
Jetson Nano Tegra X1 128 autorally 1.396 56±0.613 235uncertain1.396560.6132351.396\,56$\pm$ 0.613\,235start_ARG 1.396 56 end_ARG ± start_ARG 0.613 235 end_ARG Jetson Nano Tegra X1 128 MPPI-Generic 1.024 39±0.091 405 8uncertain1.024390.09140581.024\,39$\pm$ 0.091\,405\,8start_ARG 1.024 39 end_ARG ± start_ARG 0.091 405 8 end_ARG
Jetson Nano Tegra X1 256 autorally 1.506 45±0.575 889uncertain1.506450.5758891.506\,45$\pm$ 0.575\,889start_ARG 1.506 45 end_ARG ± start_ARG 0.575 889 end_ARG Jetson Nano Tegra X1 256 MPPI-Generic 1.383 92±0.061 362 6uncertain1.383920.06136261.383\,92$\pm$ 0.061\,362\,6start_ARG 1.383 92 end_ARG ± start_ARG 0.061 362 6 end_ARG
Jetson Nano Tegra X1 512 autorally 1.846 51±0.574 576uncertain1.846510.5745761.846\,51$\pm$ 0.574\,576start_ARG 1.846 51 end_ARG ± start_ARG 0.574 576 end_ARG Jetson Nano Tegra X1 512 MPPI-Generic 2.122 54±0.119 91uncertain2.122540.119912.122\,54$\pm$ 0.119\,91start_ARG 2.122 54 end_ARG ± start_ARG 0.119 91 end_ARG
Jetson Nano Tegra X1 1024 autorally 3.244 59±0.607 794uncertain3.244590.6077943.244\,59$\pm$ 0.607\,794start_ARG 3.244 59 end_ARG ± start_ARG 0.607 794 end_ARG Jetson Nano Tegra X1 1024 MPPI-Generic 4.105 41±1.202 12uncertain4.105411.202124.105\,41$\pm$ 1.202\,12start_ARG 4.105 41 end_ARG ± start_ARG 1.202 12 end_ARG
Jetson Nano Tegra X1 2048 autorally 5.921±0.577 712uncertain5.9210.5777125.921$\pm$ 0.577\,712start_ARG 5.921 end_ARG ± start_ARG 0.577 712 end_ARG Jetson Nano Tegra X1 2048 MPPI-Generic 7.296 89±0.398 892uncertain7.296890.3988927.296\,89$\pm$ 0.398\,892start_ARG 7.296 89 end_ARG ± start_ARG 0.398 892 end_ARG
Jetson Nano Tegra X1 4096 autorally 11.1589±0.581 031uncertain11.15890.58103111.1589$\pm$ 0.581\,031start_ARG 11.1589 end_ARG ± start_ARG 0.581 031 end_ARG Jetson Nano Tegra X1 4096 MPPI-Generic 14.1242±0.133 369uncertain14.12420.13336914.1242$\pm$ 0.133\,369start_ARG 14.1242 end_ARG ± start_ARG 0.133 369 end_ARG
Jetson Nano Tegra X1 6144 autorally 16.5458±0.605 249uncertain16.54580.60524916.5458$\pm$ 0.605\,249start_ARG 16.5458 end_ARG ± start_ARG 0.605 249 end_ARG Jetson Nano Tegra X1 6144 MPPI-Generic 21.0746±0.229 213uncertain21.07460.22921321.0746$\pm$ 0.229\,213start_ARG 21.0746 end_ARG ± start_ARG 0.229 213 end_ARG
Jetson Nano Tegra X1 8192 autorally 21.7672±0.628 532uncertain21.76720.62853221.7672$\pm$ 0.628\,532start_ARG 21.7672 end_ARG ± start_ARG 0.628 532 end_ARG Jetson Nano Tegra X1 8192 MPPI-Generic 28.0223±0.260 565uncertain28.02230.26056528.0223$\pm$ 0.260\,565start_ARG 28.0223 end_ARG ± start_ARG 0.260 565 end_ARG
Jetson Nano Tegra X1 16384 autorally 44.1579±0.947 064uncertain44.15790.94706444.1579$\pm$ 0.947\,064start_ARG 44.1579 end_ARG ± start_ARG 0.947 064 end_ARG Jetson Nano Tegra X1 16384 MPPI-Generic 56.6497±0.350 869uncertain56.64970.35086956.6497$\pm$ 0.350\,869start_ARG 56.6497 end_ARG ± start_ARG 0.350 869 end_ARG
Jetson Orin Nano (8 GB) 128 autorally 0.946 763±0.292 363uncertain0.9467630.2923630.946\,763$\pm$ 0.292\,363start_ARG 0.946 763 end_ARG ± start_ARG 0.292 363 end_ARG Jetson Orin Nano (8 GB) 128 MPPI-Generic 0.629 243±0.1296uncertain0.6292430.12960.629\,243$\pm$ 0.1296start_ARG 0.629 243 end_ARG ± start_ARG 0.1296 end_ARG
Jetson Orin Nano (8 GB) 256 autorally 0.903 75±0.066 688 6uncertain0.903750.06668860.903\,75$\pm$ 0.066\,688\,6start_ARG 0.903 75 end_ARG ± start_ARG 0.066 688 6 end_ARG Jetson Orin Nano (8 GB) 256 MPPI-Generic 0.607 396±0.022 103 4uncertain0.6073960.02210340.607\,396$\pm$ 0.022\,103\,4start_ARG 0.607 396 end_ARG ± start_ARG 0.022 103 4 end_ARG
Jetson Orin Nano (8 GB) 512 autorally 0.909 514±0.033 216 3uncertain0.9095140.03321630.909\,514$\pm$ 0.033\,216\,3start_ARG 0.909 514 end_ARG ± start_ARG 0.033 216 3 end_ARG Jetson Orin Nano (8 GB) 512 MPPI-Generic 0.728 221±0.030 154 5uncertain0.7282210.03015450.728\,221$\pm$ 0.030\,154\,5start_ARG 0.728 221 end_ARG ± start_ARG 0.030 154 5 end_ARG
Jetson Orin Nano (8 GB) 1024 autorally 1.031 62±0.030 783 8uncertain1.031620.03078381.031\,62$\pm$ 0.030\,783\,8start_ARG 1.031 62 end_ARG ± start_ARG 0.030 783 8 end_ARG Jetson Orin Nano (8 GB) 1024 MPPI-Generic 1.030 02±0.034 969 4uncertain1.030020.03496941.030\,02$\pm$ 0.034\,969\,4start_ARG 1.030 02 end_ARG ± start_ARG 0.034 969 4 end_ARG
Jetson Orin Nano (8 GB) 2048 autorally 1.408 94±0.055 614 6uncertain1.408940.05561461.408\,94$\pm$ 0.055\,614\,6start_ARG 1.408 94 end_ARG ± start_ARG 0.055 614 6 end_ARG Jetson Orin Nano (8 GB) 2048 MPPI-Generic 1.645 41±0.108 883uncertain1.645410.1088831.645\,41$\pm$ 0.108\,883start_ARG 1.645 41 end_ARG ± start_ARG 0.108 883 end_ARG
Jetson Orin Nano (8 GB) 4096 autorally 2.712 73±0.124 459uncertain2.712730.1244592.712\,73$\pm$ 0.124\,459start_ARG 2.712 73 end_ARG ± start_ARG 0.124 459 end_ARG Jetson Orin Nano (8 GB) 4096 MPPI-Generic 3.092 44±0.079 106 5uncertain3.092440.07910653.092\,44$\pm$ 0.079\,106\,5start_ARG 3.092 44 end_ARG ± start_ARG 0.079 106 5 end_ARG
Jetson Orin Nano (8 GB) 6144 autorally 3.862 98±0.100 405uncertain3.862980.1004053.862\,98$\pm$ 0.100\,405start_ARG 3.862 98 end_ARG ± start_ARG 0.100 405 end_ARG Jetson Orin Nano (8 GB) 6144 MPPI-Generic 4.545 14±0.135 722uncertain4.545140.1357224.545\,14$\pm$ 0.135\,722start_ARG 4.545 14 end_ARG ± start_ARG 0.135 722 end_ARG
Jetson Orin Nano (8 GB) 8192 autorally 5.255 01±0.131 005uncertain5.255010.1310055.255\,01$\pm$ 0.131\,005start_ARG 5.255 01 end_ARG ± start_ARG 0.131 005 end_ARG Jetson Orin Nano (8 GB) 8192 MPPI-Generic 6.102 01±0.153 504uncertain6.102010.1535046.102\,01$\pm$ 0.153\,504start_ARG 6.102 01 end_ARG ± start_ARG 0.153 504 end_ARG
Jetson Orin Nano (8 GB) 16384 autorally 9.940 03±0.196 983uncertain9.940030.1969839.940\,03$\pm$ 0.196\,983start_ARG 9.940 03 end_ARG ± start_ARG 0.196 983 end_ARG Jetson Orin Nano (8 GB) 16384 MPPI-Generic 12.0695±0.217 526uncertain12.06950.21752612.0695$\pm$ 0.217\,526start_ARG 12.0695 end_ARG ± start_ARG 0.217 526 end_ARG
CPU GPU Samples Method Avg. Time [$\mathrm{ms}$] CPU GPU Samples Method Avg. Time [$\mathrm{ms}$]
AMD 5600X 1050 Ti MPPI-Generic 128 0.214 853±0.047 800 1uncertain0.2148530.04780010.214\,853$\pm$ 0.047\,800\,1start_ARG 0.214 853 end_ARG ± start_ARG 0.047 800 1 end_ARG Intel 13900K 1050 Ti MPPI-Generic 128 0.200 793±0.059 054uncertain0.2007930.0590540.200\,793$\pm$ 0.059\,054start_ARG 0.200 793 end_ARG ± start_ARG 0.059 054 end_ARG
AMD 5600X 1050 Ti autorally 128 0.464 578±0.033 688 9uncertain0.4645780.03368890.464\,578$\pm$ 0.033\,688\,9start_ARG 0.464 578 end_ARG ± start_ARG 0.033 688 9 end_ARG Intel 13900K 1050 Ti autorally 128 0.455 909±0.159 941uncertain0.4559090.1599410.455\,909$\pm$ 0.159\,941start_ARG 0.455 909 end_ARG ± start_ARG 0.159 941 end_ARG
AMD 5600X 1050 Ti torchrl 128 46.649 427 890 777 53±2.249 891 630 328 609uncertain46.649427890777532.24989163032860946.649\,427\,890\,777\,53$\pm$ 2.249\,891\,630\,328\,609start_ARG 46.649 427 890 777 53 end_ARG ± start_ARG 2.249 891 630 328 609 end_ARG Intel 13900K 1050 Ti torchrl 128 25.603 788 137 435 888±2.713 817 147 456 561 3uncertain25.6037881374358882.713817147456561325.603\,788\,137\,435\,888$\pm$ 2.713\,817\,147\,456\,561\,3start_ARG 25.603 788 137 435 888 end_ARG ± start_ARG 2.713 817 147 456 561 3 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 256 0.226 453±0.043 313 3uncertain0.2264530.04331330.226\,453$\pm$ 0.043\,313\,3start_ARG 0.226 453 end_ARG ± start_ARG 0.043 313 3 end_ARG Intel 13900K 1050 Ti MPPI-Generic 256 0.225 711±0.057 800 7uncertain0.2257110.05780070.225\,711$\pm$ 0.057\,800\,7start_ARG 0.225 711 end_ARG ± start_ARG 0.057 800 7 end_ARG
AMD 5600X 1050 Ti autorally 256 0.464 905±0.033 803 4uncertain0.4649050.03380340.464\,905$\pm$ 0.033\,803\,4start_ARG 0.464 905 end_ARG ± start_ARG 0.033 803 4 end_ARG Intel 13900K 1050 Ti autorally 256 0.454 321±0.083 264 7uncertain0.4543210.08326470.454\,321$\pm$ 0.083\,264\,7start_ARG 0.454 321 end_ARG ± start_ARG 0.083 264 7 end_ARG
AMD 5600X 1050 Ti torchrl 256 52.884 384 155 273 416±2.402 423 955 959 606uncertain52.8843841552734162.40242395595960652.884\,384\,155\,273\,416$\pm$ 2.402\,423\,955\,959\,606start_ARG 52.884 384 155 273 416 end_ARG ± start_ARG 2.402 423 955 959 606 end_ARG Intel 13900K 1050 Ti torchrl 256 29.307 887 554 168 733±2.436 653 659 330 253 7uncertain29.3078875541687332.436653659330253729.307\,887\,554\,168\,733$\pm$ 2.436\,653\,659\,330\,253\,7start_ARG 29.307 887 554 168 733 end_ARG ± start_ARG 2.436 653 659 330 253 7 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 512 0.307 676±0.052 547uncertain0.3076760.0525470.307\,676$\pm$ 0.052\,547start_ARG 0.307 676 end_ARG ± start_ARG 0.052 547 end_ARG Intel 13900K 1050 Ti MPPI-Generic 512 0.301 609±0.065 774 3uncertain0.3016090.06577430.301\,609$\pm$ 0.065\,774\,3start_ARG 0.301 609 end_ARG ± start_ARG 0.065 774 3 end_ARG
AMD 5600X 1050 Ti autorally 512 0.481 166±0.037 496 1uncertain0.4811660.03749610.481\,166$\pm$ 0.037\,496\,1start_ARG 0.481 166 end_ARG ± start_ARG 0.037 496 1 end_ARG Intel 13900K 1050 Ti autorally 512 0.468 811±0.080 185 5uncertain0.4688110.08018550.468\,811$\pm$ 0.080\,185\,5start_ARG 0.468 811 end_ARG ± start_ARG 0.080 185 5 end_ARG
AMD 5600X 1050 Ti torchrl 512 52.965 013 504 028 35±2.365 726 109 028 937uncertain52.965013504028352.36572610902893752.965\,013\,504\,028\,35$\pm$ 2.365\,726\,109\,028\,937start_ARG 52.965 013 504 028 35 end_ARG ± start_ARG 2.365 726 109 028 937 end_ARG Intel 13900K 1050 Ti torchrl 512 29.907 815 694 808 95±2.798 864 498 586 772 3uncertain29.907815694808952.798864498586772329.907\,815\,694\,808\,95$\pm$ 2.798\,864\,498\,586\,772\,3start_ARG 29.907 815 694 808 95 end_ARG ± start_ARG 2.798 864 498 586 772 3 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 1024 0.437 319±0.054 849 6uncertain0.4373190.05484960.437\,319$\pm$ 0.054\,849\,6start_ARG 0.437 319 end_ARG ± start_ARG 0.054 849 6 end_ARG Intel 13900K 1050 Ti MPPI-Generic 1024 0.429 377±0.075 327 5uncertain0.4293770.07532750.429\,377$\pm$ 0.075\,327\,5start_ARG 0.429 377 end_ARG ± start_ARG 0.075 327 5 end_ARG
AMD 5600X 1050 Ti autorally 1024 0.519 615±0.039 718 4uncertain0.5196150.03971840.519\,615$\pm$ 0.039\,718\,4start_ARG 0.519 615 end_ARG ± start_ARG 0.039 718 4 end_ARG Intel 13900K 1050 Ti autorally 1024 0.511 98±0.096 286 5uncertain0.511980.09628650.511\,98$\pm$ 0.096\,286\,5start_ARG 0.511 98 end_ARG ± start_ARG 0.096 286 5 end_ARG
AMD 5600X 1050 Ti torchrl 1024 53.149 573 087 692 22±2.144 392 550 399 996uncertain53.149573087692222.14439255039999653.149\,573\,087\,692\,22$\pm$ 2.144\,392\,550\,399\,996start_ARG 53.149 573 087 692 22 end_ARG ± start_ARG 2.144 392 550 399 996 end_ARG Intel 13900K 1050 Ti torchrl 1024 29.902 313 709 259 01±2.995 227 000 571 960 6uncertain29.902313709259012.995227000571960629.902\,313\,709\,259\,01$\pm$ 2.995\,227\,000\,571\,960\,6start_ARG 29.902 313 709 259 01 end_ARG ± start_ARG 2.995 227 000 571 960 6 end_ARG
AMD 5600X 1050 Ti autorally 2048 0.637 062±0.043 267 9uncertain0.6370620.04326790.637\,062$\pm$ 0.043\,267\,9start_ARG 0.637 062 end_ARG ± start_ARG 0.043 267 9 end_ARG Intel 13900K 1050 Ti autorally 2048 0.629 705±0.104 016uncertain0.6297050.1040160.629\,705$\pm$ 0.104\,016start_ARG 0.629 705 end_ARG ± start_ARG 0.104 016 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 2048 0.844 269±0.061 363 3uncertain0.8442690.06136330.844\,269$\pm$ 0.061\,363\,3start_ARG 0.844 269 end_ARG ± start_ARG 0.061 363 3 end_ARG Intel 13900K 1050 Ti MPPI-Generic 2048 0.836 98±0.099 327 1uncertain0.836980.09932710.836\,98$\pm$ 0.099\,327\,1start_ARG 0.836 98 end_ARG ± start_ARG 0.099 327 1 end_ARG
AMD 5600X 1050 Ti torchrl 2048 53.575 214 624 404 97±2.453 937 143 330 288 3uncertain53.575214624404972.453937143330288353.575\,214\,624\,404\,97$\pm$ 2.453\,937\,143\,330\,288\,3start_ARG 53.575 214 624 404 97 end_ARG ± start_ARG 2.453 937 143 330 288 3 end_ARG Intel 13900K 1050 Ti torchrl 2048 30.389 081 716 537 47±2.918 950 472 504 109uncertain30.389081716537472.91895047250410930.389\,081\,716\,537\,47$\pm$ 2.918\,950\,472\,504\,109start_ARG 30.389 081 716 537 47 end_ARG ± start_ARG 2.918 950 472 504 109 end_ARG
AMD 5600X 1050 Ti autorally 4096 1.295 23±0.056 928uncertain1.295230.0569281.295\,23$\pm$ 0.056\,928start_ARG 1.295 23 end_ARG ± start_ARG 0.056 928 end_ARG Intel 13900K 1050 Ti autorally 4096 1.201 65±0.132 191uncertain1.201650.1321911.201\,65$\pm$ 0.132\,191start_ARG 1.201 65 end_ARG ± start_ARG 0.132 191 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 4096 1.483 58±0.061 568 2uncertain1.483580.06156821.483\,58$\pm$ 0.061\,568\,2start_ARG 1.483 58 end_ARG ± start_ARG 0.061 568 2 end_ARG Intel 13900K 1050 Ti MPPI-Generic 4096 1.468 12±0.134 319uncertain1.468120.1343191.468\,12$\pm$ 0.134\,319start_ARG 1.468 12 end_ARG ± start_ARG 0.134 319 end_ARG
AMD 5600X 1050 Ti torchrl 4096 55.103 763 341 903 694±2.424 383 291 043 795uncertain55.1037633419036942.42438329104379555.103\,763\,341\,903\,694$\pm$ 2.424\,383\,291\,043\,795start_ARG 55.103 763 341 903 694 end_ARG ± start_ARG 2.424 383 291 043 795 end_ARG Intel 13900K 1050 Ti torchrl 4096 31.282 362 222 671 51±2.708 110 196 085 109 5uncertain31.282362222671512.708110196085109531.282\,362\,222\,671\,51$\pm$ 2.708\,110\,196\,085\,109\,5start_ARG 31.282 362 222 671 51 end_ARG ± start_ARG 2.708 110 196 085 109 5 end_ARG
AMD 5600X 1050 Ti autorally 6144 1.459 51±0.061 768 2uncertain1.459510.06176821.459\,51$\pm$ 0.061\,768\,2start_ARG 1.459 51 end_ARG ± start_ARG 0.061 768 2 end_ARG Intel 13900K 1050 Ti autorally 6144 1.7757±0.150 588uncertain1.77570.1505881.7757$\pm$ 0.150\,588start_ARG 1.7757 end_ARG ± start_ARG 0.150 588 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 6144 2.1196±0.072 657 2uncertain2.11960.07265722.1196$\pm$ 0.072\,657\,2start_ARG 2.1196 end_ARG ± start_ARG 0.072 657 2 end_ARG Intel 13900K 1050 Ti MPPI-Generic 6144 2.103 06±0.155 404uncertain2.103060.1554042.103\,06$\pm$ 0.155\,404start_ARG 2.103 06 end_ARG ± start_ARG 0.155 404 end_ARG
AMD 5600X 1050 Ti torchrl 6144 57.429 688 930 511 47±2.136 044 589 193 735 6uncertain57.429688930511472.136044589193735657.429\,688\,930\,511\,47$\pm$ 2.136\,044\,589\,193\,735\,6start_ARG 57.429 688 930 511 47 end_ARG ± start_ARG 2.136 044 589 193 735 6 end_ARG Intel 13900K 1050 Ti torchrl 6144 34.261 162 757 873 51±2.754 586 372 387 543 7uncertain34.261162757873512.754586372387543734.261\,162\,757\,873\,51$\pm$ 2.754\,586\,372\,387\,543\,7start_ARG 34.261 162 757 873 51 end_ARG ± start_ARG 2.754 586 372 387 543 7 end_ARG
AMD 5600X 1050 Ti autorally 8192 2.076 23±0.073 865 7uncertain2.076230.07386572.076\,23$\pm$ 0.073\,865\,7start_ARG 2.076 23 end_ARG ± start_ARG 0.073 865 7 end_ARG Intel 13900K 1050 Ti autorally 8192 2.360 26±0.169 856uncertain2.360260.1698562.360\,26$\pm$ 0.169\,856start_ARG 2.360 26 end_ARG ± start_ARG 0.169 856 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 8192 2.906 15±0.084 103uncertain2.906150.0841032.906\,15$\pm$ 0.084\,103start_ARG 2.906 15 end_ARG ± start_ARG 0.084 103 end_ARG Intel 13900K 1050 Ti MPPI-Generic 8192 2.893 17±0.247 637uncertain2.893170.2476372.893\,17$\pm$ 0.247\,637start_ARG 2.893 17 end_ARG ± start_ARG 0.247 637 end_ARG
AMD 5600X 1050 Ti torchrl 8192 60.502 972 364 425 666±2.448 291 157 722 627 6uncertain60.5029723644256662.448291157722627660.502\,972\,364\,425\,666$\pm$ 2.448\,291\,157\,722\,627\,6start_ARG 60.502 972 364 425 666 end_ARG ± start_ARG 2.448 291 157 722 627 6 end_ARG Intel 13900K 1050 Ti torchrl 8192 37.079 464 197 158 8±2.684 802 959 376 344uncertain37.07946419715882.68480295937634437.079\,464\,197\,158\,8$\pm$ 2.684\,802\,959\,376\,344start_ARG 37.079 464 197 158 8 end_ARG ± start_ARG 2.684 802 959 376 344 end_ARG
AMD 5600X 1050 Ti autorally 16384 4.107 61±0.112 963uncertain4.107610.1129634.107\,61$\pm$ 0.112\,963start_ARG 4.107 61 end_ARG ± start_ARG 0.112 963 end_ARG Intel 13900K 1050 Ti autorally 16384 4.610 68±0.254 627uncertain4.610680.2546274.610\,68$\pm$ 0.254\,627start_ARG 4.610 68 end_ARG ± start_ARG 0.254 627 end_ARG
AMD 5600X 1050 Ti MPPI-Generic 16384 5.665 18±0.126 393uncertain5.665180.1263935.665\,18$\pm$ 0.126\,393start_ARG 5.665 18 end_ARG ± start_ARG 0.126 393 end_ARG Intel 13900K 1050 Ti MPPI-Generic 16384 5.637 98±0.418 242uncertain5.637980.4182425.637\,98$\pm$ 0.418\,242start_ARG 5.637 98 end_ARG ± start_ARG 0.418 242 end_ARG
AMD 5600X 1050 Ti torchrl 16384 73.403 141 975 402 87±1.930 647 192 179 201 9uncertain73.403141975402871.930647192179201973.403\,141\,975\,402\,87$\pm$ 1.930\,647\,192\,179\,201\,9start_ARG 73.403 141 975 402 87 end_ARG ± start_ARG 1.930 647 192 179 201 9 end_ARG Intel 13900K 1050 Ti torchrl 16384 51.973 490 238 189 676±2.884 466 781 077 533uncertain51.9734902381896762.88446678107753351.973\,490\,238\,189\,676$\pm$ 2.884\,466\,781\,077\,533start_ARG 51.973 490 238 189 676 end_ARG ± start_ARG 2.884 466 781 077 533 end_ARG
AMD 5600X 1650 MPPI-Generic 128 0.145 933±0.020 777 5uncertain0.1459330.02077750.145\,933$\pm$ 0.020\,777\,5start_ARG 0.145 933 end_ARG ± start_ARG 0.020 777 5 end_ARG Intel 13900K 1650 MPPI-Generic 128 0.135 448±0.026 331 9uncertain0.1354480.02633190.135\,448$\pm$ 0.026\,331\,9start_ARG 0.135 448 end_ARG ± start_ARG 0.026 331 9 end_ARG
AMD 5600X 1650 autorally 128 0.255 503±0.017 641 6uncertain0.2555030.01764160.255\,503$\pm$ 0.017\,641\,6start_ARG 0.255 503 end_ARG ± start_ARG 0.017 641 6 end_ARG Intel 13900K 1650 autorally 128 0.244 764±0.110 752uncertain0.2447640.1107520.244\,764$\pm$ 0.110\,752start_ARG 0.244 764 end_ARG ± start_ARG 0.110 752 end_ARG
AMD 5600X 1650 torchrl 128 46.523 932 695 388 815±1.960 685 391 249 216 4uncertain46.5239326953888151.960685391249216446.523\,932\,695\,388\,815$\pm$ 1.960\,685\,391\,249\,216\,4start_ARG 46.523 932 695 388 815 end_ARG ± start_ARG 1.960 685 391 249 216 4 end_ARG Intel 13900K 1650 torchrl 128 24.704 576 730 728 12±1.769 540 793 032 105 7uncertain24.704576730728121.769540793032105724.704\,576\,730\,728\,12$\pm$ 1.769\,540\,793\,032\,105\,7start_ARG 24.704 576 730 728 12 end_ARG ± start_ARG 1.769 540 793 032 105 7 end_ARG
AMD 5600X 1650 MPPI-Generic 256 0.152 956±0.018 835 4uncertain0.1529560.01883540.152\,956$\pm$ 0.018\,835\,4start_ARG 0.152 956 end_ARG ± start_ARG 0.018 835 4 end_ARG Intel 13900K 1650 MPPI-Generic 256 0.152 638±0.026 009 9uncertain0.1526380.02600990.152\,638$\pm$ 0.026\,009\,9start_ARG 0.152 638 end_ARG ± start_ARG 0.026 009 9 end_ARG
AMD 5600X 1650 autorally 256 0.256 436±0.023 154uncertain0.2564360.0231540.256\,436$\pm$ 0.023\,154start_ARG 0.256 436 end_ARG ± start_ARG 0.023 154 end_ARG Intel 13900K 1650 autorally 256 0.238 484±0.028 414 8uncertain0.2384840.02841480.238\,484$\pm$ 0.028\,414\,8start_ARG 0.238 484 end_ARG ± start_ARG 0.028 414 8 end_ARG
AMD 5600X 1650 torchrl 256 52.687 264 442 443 81±2.313 450 177 948 185uncertain52.687264442443812.31345017794818552.687\,264\,442\,443\,81$\pm$ 2.313\,450\,177\,948\,185start_ARG 52.687 264 442 443 81 end_ARG ± start_ARG 2.313 450 177 948 185 end_ARG Intel 13900K 1650 torchrl 256 28.543 481 349 945 065±1.518 854 203 689 118 2uncertain28.5434813499450651.518854203689118228.543\,481\,349\,945\,065$\pm$ 1.518\,854\,203\,689\,118\,2start_ARG 28.543 481 349 945 065 end_ARG ± start_ARG 1.518 854 203 689 118 2 end_ARG
AMD 5600X 1650 MPPI-Generic 512 0.201 461±0.021 816 4uncertain0.2014610.02181640.201\,461$\pm$ 0.021\,816\,4start_ARG 0.201 461 end_ARG ± start_ARG 0.021 816 4 end_ARG Intel 13900K 1650 MPPI-Generic 512 0.196 435±0.032 044 4uncertain0.1964350.03204440.196\,435$\pm$ 0.032\,044\,4start_ARG 0.196 435 end_ARG ± start_ARG 0.032 044 4 end_ARG
AMD 5600X 1650 autorally 512 0.264 917±0.022 641 9uncertain0.2649170.02264190.264\,917$\pm$ 0.022\,641\,9start_ARG 0.264 917 end_ARG ± start_ARG 0.022 641 9 end_ARG Intel 13900K 1650 autorally 512 0.238 714±0.024 900 6uncertain0.2387140.02490060.238\,714$\pm$ 0.024\,900\,6start_ARG 0.238 714 end_ARG ± start_ARG 0.024 900 6 end_ARG
AMD 5600X 1650 torchrl 512 52.196 183 681 488 016±2.124 205 141 479 847uncertain52.1961836814880162.12420514147984752.196\,183\,681\,488\,016$\pm$ 2.124\,205\,141\,479\,847start_ARG 52.196 183 681 488 016 end_ARG ± start_ARG 2.124 205 141 479 847 end_ARG Intel 13900K 1650 torchrl 512 28.674 972 057 342 522±1.657 040 122 220 694 3uncertain28.6749720573425221.657040122220694328.674\,972\,057\,342\,522$\pm$ 1.657\,040\,122\,220\,694\,3start_ARG 28.674 972 057 342 522 end_ARG ± start_ARG 1.657 040 122 220 694 3 end_ARG
AMD 5600X 1650 MPPI-Generic 1024 0.284 75±0.029 307 8uncertain0.284750.02930780.284\,75$\pm$ 0.029\,307\,8start_ARG 0.284 75 end_ARG ± start_ARG 0.029 307 8 end_ARG Intel 13900K 1650 MPPI-Generic 1024 0.286 556±0.034 644 5uncertain0.2865560.03464450.286\,556$\pm$ 0.034\,644\,5start_ARG 0.286 556 end_ARG ± start_ARG 0.034 644 5 end_ARG
AMD 5600X 1650 autorally 1024 0.325 993±0.022 464 2uncertain0.3259930.02246420.325\,993$\pm$ 0.022\,464\,2start_ARG 0.325 993 end_ARG ± start_ARG 0.022 464 2 end_ARG Intel 13900K 1650 autorally 1024 0.305 997±0.030 602 2uncertain0.3059970.03060220.305\,997$\pm$ 0.030\,602\,2start_ARG 0.305 997 end_ARG ± start_ARG 0.030 602 2 end_ARG
AMD 5600X 1650 torchrl 1024 52.329 415 798 187 24±1.777 293 856 785 129 5uncertain52.329415798187241.777293856785129552.329\,415\,798\,187\,24$\pm$ 1.777\,293\,856\,785\,129\,5start_ARG 52.329 415 798 187 24 end_ARG ± start_ARG 1.777 293 856 785 129 5 end_ARG Intel 13900K 1650 torchrl 1024 28.783 922 195 434 577±1.654 554 262 478 057 1uncertain28.7839221954345771.654554262478057128.783\,922\,195\,434\,577$\pm$ 1.654\,554\,262\,478\,057\,1start_ARG 28.783 922 195 434 577 end_ARG ± start_ARG 1.654 554 262 478 057 1 end_ARG
AMD 5600X 1650 autorally 2048 0.424 721±0.025 105 8uncertain0.4247210.02510580.424\,721$\pm$ 0.025\,105\,8start_ARG 0.424 721 end_ARG ± start_ARG 0.025 105 8 end_ARG Intel 13900K 1650 autorally 2048 0.404 684±0.038 143 9uncertain0.4046840.03814390.404\,684$\pm$ 0.038\,143\,9start_ARG 0.404 684 end_ARG ± start_ARG 0.038 143 9 end_ARG
AMD 5600X 1650 MPPI-Generic 2048 0.504 621±0.029 476 7uncertain0.5046210.02947670.504\,621$\pm$ 0.029\,476\,7start_ARG 0.504 621 end_ARG ± start_ARG 0.029 476 7 end_ARG Intel 13900K 1650 MPPI-Generic 2048 0.483 436±0.043 101 2uncertain0.4834360.04310120.483\,436$\pm$ 0.043\,101\,2start_ARG 0.483 436 end_ARG ± start_ARG 0.043 101 2 end_ARG
AMD 5600X 1650 torchrl 2048 52.578 122 138 977 05±2.109 717 144 827 896uncertain52.578122138977052.10971714482789652.578\,122\,138\,977\,05$\pm$ 2.109\,717\,144\,827\,896start_ARG 52.578 122 138 977 05 end_ARG ± start_ARG 2.109 717 144 827 896 end_ARG Intel 13900K 1650 torchrl 2048 29.134 824 991 226 214±1.458 942 907 895 041 3uncertain29.1348249912262141.458942907895041329.134\,824\,991\,226\,214$\pm$ 1.458\,942\,907\,895\,041\,3start_ARG 29.134 824 991 226 214 end_ARG ± start_ARG 1.458 942 907 895 041 3 end_ARG
AMD 5600X 1650 autorally 4096 0.806 484±0.044 721uncertain0.8064840.0447210.806\,484$\pm$ 0.044\,721start_ARG 0.806 484 end_ARG ± start_ARG 0.044 721 end_ARG Intel 13900K 1650 autorally 4096 0.792 676±0.061 939 6uncertain0.7926760.06193960.792\,676$\pm$ 0.061\,939\,6start_ARG 0.792 676 end_ARG ± start_ARG 0.061 939 6 end_ARG
AMD 5600X 1650 MPPI-Generic 4096 1.0046±0.044 751 5uncertain1.00460.04475151.0046$\pm$ 0.044\,751\,5start_ARG 1.0046 end_ARG ± start_ARG 0.044 751 5 end_ARG Intel 13900K 1650 MPPI-Generic 4096 0.968 006±0.063 612 3uncertain0.9680060.06361230.968\,006$\pm$ 0.063\,612\,3start_ARG 0.968 006 end_ARG ± start_ARG 0.063 612 3 end_ARG
AMD 5600X 1650 torchrl 4096 53.691 801 309 585 55±2.080 985 489 081 914uncertain53.691801309585552.08098548908191453.691\,801\,309\,585\,55$\pm$ 2.080\,985\,489\,081\,914start_ARG 53.691 801 309 585 55 end_ARG ± start_ARG 2.080 985 489 081 914 end_ARG Intel 13900K 1650 torchrl 4096 30.612 709 522 247 307±1.696 327 556 929 044 8uncertain30.6127095222473071.696327556929044830.612\,709\,522\,247\,307$\pm$ 1.696\,327\,556\,929\,044\,8start_ARG 30.612 709 522 247 307 end_ARG ± start_ARG 1.696 327 556 929 044 8 end_ARG
AMD 5600X 1650 autorally 6144 0.984 068±0.040 074 7uncertain0.9840680.04007470.984\,068$\pm$ 0.040\,074\,7start_ARG 0.984 068 end_ARG ± start_ARG 0.040 074 7 end_ARG Intel 13900K 1650 autorally 6144 0.941 817±0.062 302 2uncertain0.9418170.06230220.941\,817$\pm$ 0.062\,302\,2start_ARG 0.941 817 end_ARG ± start_ARG 0.062 302 2 end_ARG
AMD 5600X 1650 MPPI-Generic 6144 1.229 94±0.043 720 8uncertain1.229940.04372081.229\,94$\pm$ 0.043\,720\,8start_ARG 1.229 94 end_ARG ± start_ARG 0.043 720 8 end_ARG Intel 13900K 1650 MPPI-Generic 6144 1.1777±0.063 739 9uncertain1.17770.06373991.1777$\pm$ 0.063\,739\,9start_ARG 1.1777 end_ARG ± start_ARG 0.063 739 9 end_ARG
AMD 5600X 1650 torchrl 6144 55.526 461 601 257 37±1.795 973 822 927 248 2uncertain55.526461601257371.795973822927248255.526\,461\,601\,257\,37$\pm$ 1.795\,973\,822\,927\,248\,2start_ARG 55.526 461 601 257 37 end_ARG ± start_ARG 1.795 973 822 927 248 2 end_ARG Intel 13900K 1650 torchrl 6144 32.394 740 343 093 86±1.724 235 578 885 225 5uncertain32.394740343093861.724235578885225532.394\,740\,343\,093\,86$\pm$ 1.724\,235\,578\,885\,225\,5start_ARG 32.394 740 343 093 86 end_ARG ± start_ARG 1.724 235 578 885 225 5 end_ARG
AMD 5600X 1650 autorally 8192 1.386 96±0.047 558 7uncertain1.386960.04755871.386\,96$\pm$ 0.047\,558\,7start_ARG 1.386 96 end_ARG ± start_ARG 0.047 558 7 end_ARG Intel 13900K 1650 autorally 8192 1.334 33±0.072 107 9uncertain1.334330.07210791.334\,33$\pm$ 0.072\,107\,9start_ARG 1.334 33 end_ARG ± start_ARG 0.072 107 9 end_ARG
AMD 5600X 1650 MPPI-Generic 8192 1.774 03±0.051 467 2uncertain1.774030.05146721.774\,03$\pm$ 0.051\,467\,2start_ARG 1.774 03 end_ARG ± start_ARG 0.051 467 2 end_ARG Intel 13900K 1650 MPPI-Generic 8192 1.678 93±0.075 795 8uncertain1.678930.07579581.678\,93$\pm$ 0.075\,795\,8start_ARG 1.678 93 end_ARG ± start_ARG 0.075 795 8 end_ARG
AMD 5600X 1650 torchrl 8192 57.844 109 296 798 77±2.053 263 522 175 228uncertain57.844109296798772.05326352217522857.844\,109\,296\,798\,77$\pm$ 2.053\,263\,522\,175\,228start_ARG 57.844 109 296 798 77 end_ARG ± start_ARG 2.053 263 522 175 228 end_ARG Intel 13900K 1650 torchrl 8192 34.515 250 444 412 175±1.458 090 726 785 84uncertain34.5152504444121751.4580907267858434.515\,250\,444\,412\,175$\pm$ 1.458\,090\,726\,785\,84start_ARG 34.515 250 444 412 175 end_ARG ± start_ARG 1.458 090 726 785 84 end_ARG
AMD 5600X 1650 autorally 16384 2.703 91±0.068 463 3uncertain2.703910.06846332.703\,91$\pm$ 0.068\,463\,3start_ARG 2.703 91 end_ARG ± start_ARG 0.068 463 3 end_ARG Intel 13900K 1650 autorally 16384 2.590 31±0.099 942 4uncertain2.590310.09994242.590\,31$\pm$ 0.099\,942\,4start_ARG 2.590 31 end_ARG ± start_ARG 0.099 942 4 end_ARG
AMD 5600X 1650 MPPI-Generic 16384 3.293 37±0.110 754uncertain3.293370.1107543.293\,37$\pm$ 0.110\,754start_ARG 3.293 37 end_ARG ± start_ARG 0.110 754 end_ARG Intel 13900K 1650 MPPI-Generic 16384 3.128 43±0.127 66uncertain3.128430.127663.128\,43$\pm$ 0.127\,66start_ARG 3.128 43 end_ARG ± start_ARG 0.127 66 end_ARG
AMD 5600X 1650 torchrl 16384 68.918 861 150 741 49±1.734 290 582 198 782 6uncertain68.918861150741491.734290582198782668.918\,861\,150\,741\,49$\pm$ 1.734\,290\,582\,198\,782\,6start_ARG 68.918 861 150 741 49 end_ARG ± start_ARG 1.734 290 582 198 782 6 end_ARG Intel 13900K 1650 torchrl 16384 46.419 321 298 599 22±1.730 473 429 964 794uncertain46.419321298599221.73047342996479446.419\,321\,298\,599\,22$\pm$ 1.730\,473\,429\,964\,794start_ARG 46.419 321 298 599 22 end_ARG ± start_ARG 1.730 473 429 964 794 end_ARG